### v1

# TinyMyo: a Tiny Foundation Model for Flexible EMG Signal Processing at the Edge

Abstract
Surface electromyography (EMG) is a non-invasive sensing modality used in several domains, including biomechanics, rehabilitation, prosthetic control, and emerging human-machine interaction paradigms. Despite decades of use, significant challenges remain in achieving robust generalization across subjects, recording systems, and acquisition protocols. To tackle these challenges, foundation models (FMs) are gaining traction when targeting end-to-end applications based on EMG signals. Yet, existing EMG FMs remain limited to single downstream tasks and lack deployability on embedded platforms.
In this work, we present TinyMyo, a lightweight FM based on a Transformer encoder architecture. The model is pre-trained in a self-supervised manner on publicly available datasets and achieves high reconstruction fidelity with only 3.6M parameters.
With minimal task-specific head adaptations, the same backbone is used to tackle multiple downstream tasks, leveraging datasets acquired from diverse sensing locations and hardware platforms. We demonstrate generalization across hand gesture classification, hand kinematic regression, speech production and recognition, with performance comparable to or surpassing the state of the art (SoA), and model size below 5M parameters. We achieve SoA results compared to previous FM-based works on the NinaPro DB5 ($89.4\pm 0.16\%$), UCI-EMG ($97.56\pm 0.32\%$), and EPN-612 ($96.74\pm 0.09\%$) datasets.
We report, to the best of our knowledge, the first deployment of an EMG FM on an ultra-low-power microcontroller (GAP9), achieving an average power envelope of $36.45\text{mW}$.
By open-sourcing the pre-trained and the downstream task architectures (https://github.com/pulp-bio/BioFoundation), we aim to provide a flexible resource that can accelerate future research and serve as a common foundation for the EMG community.
{IEEEkeywords}
EMG, self-supervised, foundation model, edgeAI, gesture recognition, kinematic regression, speech
1 Introduction
Electromyography (EMG) measures variations in the electrical activity of muscles during contractions, reflecting the motor-neuron activation patterns [farina_2023_rev]. EMG can be acquired using invasive methods (needle EMG), suitable for a detailed characterization of individual motor units, or using non-invasive, surface-based techniques, to capture the global activity of muscle fibres  [farina_2023_rev, avila_2023_rev].
Surface EMG has found extensive application across a wide range of domains, including biomechanical analysis and rehabilitation [avila_2023_rev], human–robot interaction [xiong_emgrobot_2024], prosthetic control [cimolato_2022emglower, marinelli_2023emgupper], and emerging human–computer interaction interfaces [generic_neuromotor_interface_2025]. However, despite decades of widespread use, many important challenges remain.
First, EMG signals are highly affected by noise, motion artifacts, and cross-talk from adjacent muscles [boyer_emg_noise]. Also, acquisition systems present large variability in terms of number and arrangement of electrodes, sensor geometry, sampling rates, amplification stages, and filtering pipelines [merletti_2020_tutorial]. These sources of variability pose challenges to the design of models that can generalize across heterogeneous hardware configurations. In addition, EMG recordings are strongly influenced by both inter-subject and intra-subject variability [xiong_2021_review]. Differences in anatomy, subcutaneous fat, electrode placement, skin impedance, and day-to-day physiological or behavioural factors lead to substantial shifts in the EMG distribution, posing an additional challenge to generalization across users or recording sessions [ni_2024_survey].
Foundation Models (FM) offer a great promise to address these challenges. Inspired by the success of self-supervised pre-training in vision, language, and speech, as well as by emerging applications in biosignals [gu_2025_foundationsurvery], the core hypothesis is that large-scale datasets can be leveraged to learn high-level, platform and subject-agnostic signal representations. For EMG, such models could provide a unified backbone for many downstream tasks, building upon a robust, device-agnostic feature extraction, improving generalization across devices and subjects. However, progress in FMs of EMG has been limited by the scarcity of large sEMG corpora, and only with recent large-scale datasets (collected from thousands of users) [salter_emg2pose_2024] the exploration of generalized FMs of EMG became viable.
In this context, existing EMG FMs present significant limitations. Current models mostly focus on a single task (typically gesture classification) [chen_2025_physiowave, chen_2025_waveformer]. Furthermore, the size of EMG FMs remains large, and no prior work demonstrated the deployment of an EMG FM on embedded hardware. This motivates the need for efficient FM architectures that meet the memory and computational constraints of embedded devices [benatti_2025_emgreview].
We address these limitations by proposing the following contribution.
  • We present TinyMyo, a lightweight Transformer-based encoder operating directly on EMG signals in the time domain. TinyMyo is pre-trained using a self-supervised masked-reconstruction framework on publicly available, heterogeneous sEMG datasets, achieving high reconstruction fidelity with only 3.6 million parameters.
  • Unlike previous EMG FMs, which are typically designed around a single downstream task or rely on large, domain-agnostic encoders, TinyMyo generalizes across heterogeneous sEMG datasets and sensor configurations. Our architecture follows the general Transformer design adopted in prior EMG works [chen_2025_physiowave, chen_2025_waveformer], formulated in a unified EMG FM able to support a wide range of downstream tasks using the same pretrained backbone. We demonstrate the versatility of the pre-trained encoder through minimal task-specific head adaptations across diverse downstream tasks, including hand-gesture classification on the Ninapro DB5 [ninapro_db5], EPN-612 [benalcazar_emg-epn-612_2020], UC Irvine [emg_data_for_gestures_481] and Meta’s Generic Neuromotor Interface [generic_neuromotor_interface_2025] datasets, hand-kinematic regression on the Ninapro DB8 [ninapro_db8], and speech production and discrimination on the Gaddy Silent Speech Dataset [gaddy-klein-2020-digital]. Our model achieves competitive performances across all downstream tasks, either surpassing SoA or guaranteeing a comparable accuracy with a sub-5M class model size.
  • We demonstrate for the first time the deployment of an EMG FM on an ultra-low-power microcontroller (GAP9), achieving an average power envelope of $36.45\text{mW}$.
  • We open source the model and weights to enable reproducibility and further research by the community.
2 Related Works
Several efforts have been made to develop FMs for biosignals, and most of the works focused on electrocardiography (ECG), photoplethysmography (PPG), and electroencephalography (EEG) [ni_2024_survey]. In contrast, research dedicated to EMG FMs remains limited.
Turgut et al. [turgut_2024_otis] introduced OTiS, a foundation model with 45 M parameters pre-trained on a broad collection of modalities, including ECG, temperature, audio, electromechanical signals, EEG, and financial time series. Their work demonstrated the feasibility of handling heterogeneous time-series data, including diverse input types, sampling rates, and channel configurations, within a unified pre-training framework. In the original study, EMG is used exclusively for multi-class muscular disease classification. Subsequent works by Cheng et al. [chen_2025_physiowave, chen_2025_waveformer] evaluated the same model on hand-gesture classification using the EPN612 [benalcazar_emg-epn-612_2020], NinaPro DB5 [ninapro_db5], NinaPro DB6 [ninapro_db6], and UCI EMG [emg_data_for_gestures_481] datasets, demonstrating high accuracy across all datasets.
Goswami et al. proposed MOMENT [goswami_2024_moment], a 350-M parameter FM pre-trained on a wide range of domains and temporal resolutions, including electrical signals, traffic, weather, disease progression, motion capture, simulated data, and ECG (among biopotentials). In the original work, EMG is not included in the pre-training corpus nor in the downstream evaluations. However, recent studies by Cheng et al. [chen_2025_waveformer, chen_2025_physiowave] assessed MOMENT on the hand-gesture classification task using the EPN612, NinaPro DB5, NinaPro DB6, and UCI EMG datasets, demonstrating the applicability of MOMENT as an EMG FM.
PhysioWave [chen_2025_physiowave] introduced a modality-specific FM for EMG and ECG. The authors further propose a unified multimodal framework combining these encoders with existing pre-trained EEG models, leveraging a multi-scale wavelet-transformer architecture, using wavelet decomposition to capture time-frequency features at different scales. For the EMG-specific model, the primary downstream task is hand-gesture classification, evaluated on NinaPro DB5, EPN612, and UCI EMG. PhysioWave achieved a performance comparable to Moment and OTIS with a much smaller parameter count (5M for the smaller model).
The same authors proposed WaveFormer [chen_2025_waveformer], a lightweight Transformer model tailored for sEMG-based gesture recognition, with a model size of only 3.1 M parameters. To ensure the extraction of frequency-domain features remains compact, the model integrates a learnable wavelet transform based on multi-level wavelet decomposition and depthwise separable convolutions. Downstream evaluations are conducted on the EPN612, NinaPro DB5, NinaPro DB6, and UCI EMG datasets, further improving the performance over PhysioWave.
More recently, Jiang et al. [jiang_2025_towards] introduced PhysioOmni, a multi-modal FM designed to handle a wide range of physiological signals such as EEG, ECG, EOG and EMG. For each modality, the authors use the same encoder as in LaBraM [jiang_2024_labram] (5.85.8M parameters per modality, considering its smallest-Base version). They adopt a decoupled multimodal tokenizer and masked signal modelling to ensure robustness under arbitrary missing modalities during inference. During pre-training, PhysioOmni relies on a fixed set of modalities (EEG, EOG, and EMG), which makes the architecture unsuitable for pre-training when one modality is missing. For downstream evaluation, the authors include 5-class sleep stage classification on the HMC [alvarez_2021_hmc] dataset and gait kinematic regression on the FBM [brantley_2018_fbm] dataset.
In this context, existing EMG FMs appear limited by the small number of explored downstream tasks, as well as by the lack of deployability on embedded platforms. TinyMyo is the first EMG FM to cover a broader range of downstream tasks with SoA performance, as will be extensively described in Section 4, while also being deployed on an ultra-low-power MCU.
3 Methods
Figure 1: Overview of the TinyMyo architecture. (a) Pre-training framework: the input signal is first tokenized using a channel-independent patching strategy with random masking, processed by eight bidirectional Transformer Encoder blocks, and reconstructed by a Lightweight Decoder. (b) Downstream architectures: after pre-training, the Decoder is removed and replaced with task-specific heads. For Hand-Gesture Classification and Kinematic Regression, a linear head is used to produce output classes or joint angles. For speech-related tasks, the EMG input is first downsampled through residual convolutional blocks and then passed through the pre-trained Transformer Encoder, followed by a linear head. The output is text for the Speech Recognition task. For the the Speech Production task, a vocoder (HiFi-GAN) is added to produce audio from the predicted MFCC features. (c) Model Deployment: the architecture for the hand-gesture classification task is implemented on the GAP9 microcontroller.
Figure 1 shows the TinyMyo framework. In the following sections, we present the strategy adopted for model pre-training (Sect. 3.1) and downstream tasks (Sect. 3.2).
3.1 Foundation Model Pre-Training
3.1.1 Tokenization
TinyMyo adopts a Channel-Independent patching strategy, following methodologies effectively applied in time-series forecasting [nietime] and vision [bao2024_channel_visiontransformersimage]. Unlike other recent biosignal (EEG) foundation models that mix spatial information during the tokenization phase via 2D convolutions that span multiple channels (e.g., FEMBA uses a $2\times 16$ kernel [tegon_2025_femba]) or project them into a unified latent space (e.g., LUNA [doner2025luna]), we strictly maintain per-channel granularity. This design choice is grounded in the physiological nature of EMG recordings, where each channel corresponds to a distinct anatomical location. By treating the channels independently at the embedding stage (using a $1\times L$ kernel), the model allows the subsequent transformer layers to learn inter-channel dependencies via attention, a strategy proven effective for biosignal decoding [dimofte_2025_cerebro].
Formally, let the input EMG waveform be denoted as $\mathbf{X}\in\mathbb{R}^{T\times C}$, with $T$ timesteps and $C$ channels. We apply a sliding window with patch length $L$ and stride $S$ independently to each channel. This yields a grid of patches $\mathbf{P}\in\mathbb{R}^{C\times N_{p}\times L}$, where $N_{p}=\lfloor(T-L)/S\rfloor+1$ is the number of temporal patches per channel.
We employ a window of $T=1000$ timesteps, $L=S=20$, and $C=16$, yielding to $50$ tokens per channel ($N_{p}$) and a total sequence length of $800$ tokens ($N$).
To form the input sequence for the Transformer, we flatten the channel and temporal dimensions, resulting in a total sequence length of $N=C\times N_{p}$. Each patch $\mathbf{P}_{c,i}$ (where $c$ is the channel index and $i$ is the patch index) is mapped to a latent embedding of dimension $d_{e}$ via a shared learnable linear projection $\mathbf{W}_{\mathrm{proj}}\in\mathbb{R}^{d_{e}\times L}$. The resulting token sequence $\mathbf{E}\in\mathbb{R}^{N\times d_{e}}$ consists of embeddings:
$\mathbf{E}_{k}=\mathbf{W}_{\mathrm{proj}}\,\mathbf{P}_{c,i}^{\top},$ (1)
where the token index $k$ corresponds to the flattened arrangement of channel-time pairs $(c,i)$. We do not add learnable positional embeddings; instead, we inject positional information via Rotary Position Embeddings (RoPE) [su_2023_roformerenhancedtransformerrotary] within the attention blocks.
3.1.2 Architecture
We adopt an Encoder-Decoder architecture.
The Encoder (see Fig. 1) consists of 8 identical layers, based on pre-LayerNorm Transformer encoder [vaswani_2023_attentionneed, xiong_2020_layernormalizationtransformerarchitecture], with 3 attention heads and an embedding dimension of 192, analogous to the configuration used in EfficientViT [liu2023efficientvitmemoryefficientvision].
We adopt a RoPE approach, which introduces position information by rotating query and key vectors in 2D planes instead of adding a position vector; such a method has already been adopted for biosignals [wang_2024_eegpt, chen_2025_physiowave, chen_2025_waveformer].
During pre-training, the Encoder’s output is passed to a lightweight Decoder (3.9k parameters), consisting of a linear layer that outputs a sequence $\mathbf{\hat{P}}$, which is a reconstruction of the original patch sequence $\mathbf{P}$. Our approach draws inspiration from SimMIM [xie2022simmimsimpleframeworkmasked], which demonstrates that a minimal reconstruction head (e.g., a single linear layer) is sufficient for effective self-supervised pre-training.
The use of a lightweight decoder is intentional: placing the full reconstruction burden on the encoder forces it to learn semantically rich and task-agnostic representations, rather than relying on a high-capacity decoder to compensate for missing information. This design strategy is supported by recent survey analyses in the masked image modeling (MIM) literature: shallow decoders tend to limit shortcut learning and encourage the formation of stronger latent representations within the encoder [zhang2022surveymaskedautoencoderselfsupervised]. As a result, the learned features exhibit improved generalization across downstream tasks and recording conditions. Further details on the model’s parameters are reported in Table 1 (left).
3.1.3 Pre-training objectives
We adopt a random-masking approach [xie2022simmimsimpleframeworkmasked], where a random subset $\mathcal{M}$ of tokens is replaced by a learnable $\mathcal{[MASK]}$ token. Applying the masking over each patch enforces contextual inference over tens of milliseconds (physiologically meaningful burst segments), rather than trivial gap filling. We adopt a masking ratio of $50\%$ as in previous works [dimofte_2025_cerebro, tegon_2025_femba].
We employ the Smooth L1 Loss function as reconstruction objective between masked and original patches, as in [chen_2025_physiowave]. We define the following loss components:
$\mathcal{L}_{\mathrm{masked}}=\frac{1}{|\mathcal{MASK}|}\sum_{(c,i)\in\mathcal{MASK}}\mathrm{SmoothL1}(\mathbf{P}_{c,i},\mathbf{\hat{P}}_{c,i}),$ (2)
$\mathcal{L}_{\mathrm{visible}}=\frac{1}{|\mathcal{M}|}\sum_{(c,i)\in\mathcal{M}}\mathrm{SmoothL1}(\mathbf{P}_{c,i},\mathbf{\hat{P}}_{c,i}),$ (3)
where $\mathcal{MASK}$ and $\mathcal{M}$ are the sets of masked and visible patches, respectively. To avoid overfitting on unmasked input, the loss on visible patches is reduced by a factor of $0.1$. The final reconstruction objective is given by:
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{masked}}+0.1\cdot\mathcal{L}_{\mathrm{visible}}$ (4)
Additional details on pre-training set-up are provided in Table 1 (right).
Table 1: Hyperparameters for masked EMG pre-training.
Hyperparameter Value
Transformer Encoder
Timesteps 1000
Patch size $(H{\times}W)$ $1\times 20$
Channels 16
Embed dim 192
Layers 8
Heads 3
QKV bias Yes
QK norm No
MLP ratio 4.0
MLP size 768
Attn drop 0.1
Proj drop 0.1
Drop-path 0.1
Decoder
Decoder embed dim 192
Setup Value
Batch size 512
Peak LR $1\!\times\!10^{-4}$
Minimum LR $1\!\times\!10^{-6}$
Optimizer $(\beta_{1},\beta_{2})$ AdamW (0.9,0.98)
LR schedule Cosine
Weight decay 0.01
Epochs / warm-up 50 / 10
Grad accum batches 8
Grad clip 3
Mask ratio 0.5
Max seq length 1000
LR: learning rate
All experiments are implemented in Python 3.10 using PyTorch Lightning and Hydra for modularity, configurability, and reproducibility. The proposed foundation model is trained on the CSCS Alps HPC infrastructure using NVIDIA GH200 GPUs, employing a single node with 4x NVIDIA GH200 GPUs in Distributed Data Parallel (DDP) mode.
3.1.4 Pre-training Datasets
For model pre-training, we employ three publicly available datasets, with a summary provided in Table 2.
Table 2: Pre-training EMG datasets.
Dataset #Sub $f_{s}\text{[Hz]}$ #Ch Size [GB] Placement
Ninapro DB6 [ninapro_db6] 10 2000 14 20.3 Forearm
Ninapro DB7 [ninapro_db6] 22 2000 12 30.9 Forearm
EMG2Pose [salter_emg2pose_2024] 192 2000 16 431 Wrist
Ninapro DB6 [ninapro_db6] contains sEMG from 10 subjects performing 7 hand grasp types, each repeated 12 times over 5 separate recording days. Signals were acquired with Trigno sensors (14 channels, $2~\text{KHz}$ sampling rate), placed on the forearm. The total dataset size is $20.3~\text{GB}$.
Ninapro DB7[ninapro_db7] contains data from 20 intact subjects and 2 transradial amputees, performing 40 movements spanning isolated finger/wrist actions and grasp patterns. Data were collected using Trigno sensors (12 channels, $2~\text{kHz}$ sampling rate), placed on the forearm. The total dataset size is $30.9~\text{GB}$.
EMG2Pose[salter_emg2pose_2024] contains data from 193 participants, performing 29 different hand motions. Data were recorded with a custom wristband (16 channels, $2~\text{kHz}$ sampling rate). The total dataset size is $431~\text{GB}$.
3.1.5 Pre-training Data Processing
EMG data are pre-processed using a Butterworth band-pass filter ($4^{th}$ order, $20-450~\text{Hz}$), followed by a notch filter ($50~\text{Hz}$), as in [chen_2025_physiowave]. Min-Max, channel-wise normalization ([1,1][-1,1] range) is applied. Each recording is then segmented into fixed-length windows of 1000 samples ($500~\text{ms}$ at $2~\text{kHz}$) with 50% overlap, providing sufficient temporal context [smith_emgwindow_2011][delafuente_emgwindow_2021] while controlling computational load.
Datasets with fewer than 16 channels are zero-padded. Given that each training batch might contain data from each dataset, this approach ensures consistency in the number of channels per batch.
We did not apply any data augmentation in line with previous works [chen_2025_physiowave, chen_2025_waveformer] which already showed SoA performances. Additionally, given the small model size (3.6 M parameters) and the limited effect of data augmentation shown in scaling-laws regimes [kaplan2020scalinglawsneurallanguage], we rely on the sample-efficiency of the architecture rather than increasing artificially the data variance.
3.2 Downstream Tasks
3.2.1 Pipeline Overview
For all downstream tasks, the Encoder architecture described in Sec. 3.1.2 is retained without modification, while the Decoder used during pretraining is removed.
The Encoder output is processed through a channel-fusion and temporal pooling module (detailed in the following subsection), after which the resulting representation is fed into task-specific prediction heads, whose specifics are provided in the corresponding downstream task sections (see Sec.3.2.2).
For each temporal patch index $p$, the Encoder outputs a set of channel-wise embeddings $\{\mathbf{z}_{c,p}\in\mathbb{R}^{d_{e}}\}_{c=1}^{C}$. We fuse these representations by concatenation, producing a per-patch vector $\mathbf{z}_{1,p}\|\cdots\|\mathbf{z}_{C,p}]\in\mathbb{R}^{C\times d_{e}}$. This choice preserves electrode-specific information by retaining the per-channel structure, allowing the downstream linear head to assign independent weights to each channel. Ablation studies confirmed that concatenation consistently outperforms mean-based fusion.
[…]
3.2.2 Downstream Tasks
Table 3: Summary of downstream EMG datasets.
Dataset #Subj. $f_{s}[\text{Hz}]$ Ch Size [GB] Placement Window size (ms) Pre-Processing Metrics
Ninapro DB5 [ninapro_db5] 10 200 16 0.2 Forearm 200, 25% overlap $||$ 1000, 25% overlap BP 20–90 Hz $\rightarrow$ NC 50 Hz Accuracy, F1, AUROC
EPN-612 [benalcazar_emg-epn-612_2020] 612 200 8 5.5 Forearm 200, 25% overlap $||$ 1000, 25% overlap BP 20–90 Hz $\rightarrow$ NC 50 Hz Accuracy, F1, AUROC
UC Irvine [emg_data_for_gestures_481] 36 200 8 0.32 Forearm 200, 25% overlap $||$ 1000, 25% overlap BP 20–90 Hz $\rightarrow$ NC 50 Hz Accuracy, F1
Generic Neuromotor Interface [generic_neuromotor_interface_2025] 100 2000 16 31.1 Hand/Arm 8000 (no overlap) HP ($40\text{Hz}$) CLER
Ninapro DB8 [ninapro_db8] 12 2000 16 23.6 Forearm 200 MAE
Gaddy Silent Speech [gaddy-klein-2020-digital] 1 1000 8 3.9 Face/Neck $\approx$3000 Series of NC ($60,...,420]\text{Hz} \rightarrow$ HP $2\text{Hz}$ $\rightarrow$ downsampling WER
BP: band-pass filter; NC: notch filter.
The pretrained encoder is evaluated on the datasets listed in Table 3. We classify the downstream tasks into three categories: Hand Gesture Classification, Hand Kinematic Regression, and Speech Synthesis and Recognition.
Downstream 1: Hand Gesture Classification
This task consists of classifying different types of hand-wrist movements and is evaluated on the following datasets:
  • Ninapro DB5 [ninapro_db5] contains data collected from 10 intact subjects, using two Thalmic Myo Armbands (16 channels, $200~\text{Hz}$). The task consists of classifying 52 hand-wrist movements.
  • EPN-612 [benalcazar_emg-epn-612_2020] contains data collected from 612 subjects, using the Myo Armband (8 channels, $200~\text{Hz}$). The task consists of classifying 5 hand movements.
  • UC Irvine [emg_data_for_gestures_481] contains data from 36 subjects (8 channels, $200~\text{Hz}$). The task consists of classifying 7 movements.
Datasets are pre-processed using a band-pass filter (Butterworth, $4^{th}$ order, $20-90]~\text{Hz}$), followed by a notch filter ($50~\text{Hz}$). Then, windowing is applied. We perform an ablation over two different window sizes: $200~\text{ms}$ with 25% overlap, and $1000~\text{ms}$ with 25% overlap. After filtering and windowing, each dataset is normalized using per-channel z-score normalization.
Compared to the pre-training, datasets with fewer than 16 channels are not zero-padded because tokenization is performed on a per-channel basis; thus, having fewer channels simply results in fewer tokens, eliminating the need for padding.
For Ninapro DB5, we perform a per-repetition split as in [chen_2025_physiowave], each recording contains six repetitions, and for each subject, we assign whole repetitions to a single split so that no subject data overlaps across splits. Specifically, repetitions 1, 3, 4, and 6 are used for training, repetition 2 for validation, and repetition 5 for testing.
The UCI EMG dataset is split by subject index into three disjoint groups: subjects 1–24 are used for training, subjects 25–30 for validation, and subjects 31–36 for testing. Only subjects 11 and 30 performed the ”extended palm” gesture, so they were excluded from all data splits, hence reducing the total number of gestures to 6.
For EPN-612, we preserve the original per-user organization [benalcazar_emg-epn-612_2020] and derive splits per user. When constructing the testing split, held-out examples per user are used, ensuring that samples from the same user do not appear across train, validation, or test splits.
The model head architecture is lightweight and consists of a single fully connected linear layer that maps the pooled feature representation (max $192\times N_{c}$, $N_{c}=$ number of channels ) to the final class logits ($N_{o}$, $N_{o}=$number of outputs). No additional hidden layers, normalization modules, or non-linearities are used in the head. Hence, the model head has a maximum of $192\times N_{c}\times N_{o}$ parameters. For training, we employ the cross-entropy loss computed over the predicted logits and the corresponding ground-truth labels. We evaluate performance using Accuracy, F1-score, and AUROC, with the respective definitions given below:
Accuracy (micro) $=\frac{1}{N}\sum_{i=1}^{N}1(\hat{y_{i}}=y_{i}),$
Accuracy (macro) $=\frac{1}{K}\sum_{k=1}^{K}\frac{\mathrm{TP}_{k}+\mathrm{TN}_{k}}{\mathrm{TP}_{k}+\mathrm{TN}_{k}+\mathrm{FP}_{k}+\mathrm{FN}_{k}},$
F1-score (macro) $=\frac{1}{K}\sum_{k=1}^{K}\frac{2\,\mathrm{TP}_{k}}{2\,\mathrm{TP}_{k}+\mathrm{FP}_{k}+\mathrm{FN}_{k}},$
AUROC (macro) $=\frac{1}{K}\sum_{k=1}^{K}\mathrm{AUROC}_{k}.$
where $\hat{y}_{i}$ is the predicted class, $y_{i}$ the ground-truth label, and $\mathrm{TP}_{k}$, $\mathrm{FP}_{k}$, $\mathrm{FN}_{k}$ denote true positives, false positives, and false negatives for class $k$.
We repeat each experiment five times using five distinct random seeds and report the mean performance across runs.
An additional dataset is also considered as part of this downstream task:
  • Generic Neuromotor Interface [generic_neuromotor_interface_2025] contains data collected from 4900 participants, recorded with a custom wrist-band developed by Meta Reality Labs (16 channels, $2\text{kHz}$). Data from only 100 subjects are publicly available [metagit], randomly selected from the entire cohort.
Specifically, we tackle the discrete-gesture task presented in the original work, consisting of the classification of 9 fine finger movements. We treat this task separately, compared to the three other datasets presented above, owing to substantial methodological differences in data processing and evaluation protocols.
No pre-processing is applied to the data. Signals are windowed using $8000~\text{ms}$ windows, no overlap. The train dataset contains 80% of data from each subject for train, 10% for validation, 10% for test. The model head consists of a single linear layer, with input size of $N_{c}\times 192$, output size $9$. We employ the Cross-Entropy Loss function, and we conduct 5 runs with 5 seeds values.
We evaluate the performance with the Classification Error Rate (CLER), defined as the proportion of ground-truth labels for which the matching prediction is incorrect [generic_neuromotor_interface_2025]. The CLER is computed per-gesture and reported as the average across gestures. The CLER formula is given by:
$\mathrm{CLER}=\frac{1}{|N_{G}|}\sum_{g\in G}\left(1-\frac{C_{g,g}}{\sum_{p\in G}C_{g,p}}\right).$
where $N_{G}$ is the total number of gestures, $g\in N_{G}$, $C_{g,g}$ is the number of correct predictions for gesture $g$ and $\sum_{p\in G}C_{g,p}$ is the total number of ground-truth occurrences of $g$. In the original work, the architecture is tested on the full sequence of EMG data. The architecture proposed in this work cannot process the entire sequence due to quadratic complexity. We adopt a windowed inference approach , which involves sliding a window of $8\text{s}$, with $2\text{s}$ stride. In order to compare the original LSTM architecture with the proposed Transformer model, the same windowed inference strategy is applied to the LSTM during evaluation, and both results are reported.
Downstream 2: Hand Kinematic Regression
This task consists of predicting continuous hand and wrist movement (regression task). We employ the Ninapro DB8 [ninapro_db8] dataset. It contains EMG data collected from 12 subjects (16 channels, $2~\text{kHz}$), with 3 sessions per subject. Data were collected with an instrumented glove from ground-truth. The task consists of regressing 5 Degrees of Freedom (DoFs), defined as linear combinations of the 18 DoFs of the instrumented glove [krasoulis_2019_effect].
Each EMG channel is independently standardized using z-score normalization. No filtering is applied since the original data is already released filtered ($4\textsuperscript{th}$ order bandpass Butterworth filter, $10-500~\text{Hz}]$. The glove data undergo a mapping from 18 Degrees of Freedom (DoF) to 5 Degrees of Actuation (DoA). Both EMG and glove data are then segmented into fixed-size windows with no overlap in two different settings: $1000\text{ms}$ and $200\text{ms}$.
Following the author’s recommendations, we consider the first two sessions for training, the third as validation. The model head architecture (788 K parameters) consists of a compact series of pointwise and depthwise convolutions, followed by upsampling to a fixed length. We employ the L1 loss function.
We report the mean absolute error (MAE), root mean squared error (RMSE), and the coefficient of determination $R^{2}$, defined as:
$\mathrm{MAE}=\frac{1}{N}\sum_{i=1}^{N}\lvert y_{i}-\hat{y}_{i}\rvert$.
$\mathrm{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}.$
$R^{2}=1-\frac{\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}.$
where $y_{i}$ are the true values, $\hat{y}_{i}$ the predicted values, and $\bar{y}=\frac{1}{N}\sum_{i=1}^{N}y_{i}$. We conduct 5 runs, with 5 random seed values, and report the average.
Downstream 3: Speech Synthesis and Recognition
We employ the open-vocabulary dataset presented by Gaddy et al. [gaddy-klein-2020-digital]. It comprises 8 channels, sampled at $1000~\text{Hz}$, recorded using wet electrodes, placed on the face. Data were collected from one subject in multiple sessions while reading pieces of public domain books. The dataset contains both vocalized (i.e., normal articulation with sound production) and silent (i.e., articulation without sound production) EMG data. In the vocalized setting, EMG data were recoded together with a microphone, serving as ground truth. For the silent setting, audio recorded during vocalized experiments is aligned to the corresponding silent EMG data using the dynamic time warping algorithm [gaddy-klein-2020-digital]. The publicly available dataset contains already-aligned data.
For both tasks, we adopt the same data processing pipeline as in [gaddy2021improved]. A cascade of notch filters is applied to suppress the first six harmonics of power line frequency ($60,240,...,420]~\text{Hz}$). Additionally, a high-pass filter (Butterworth, $2~\text{Hz}$, 3rd order) is applied. Data are downsampled (from $1000~\text{Hz}$ to $516.79~\text{Hz}$, matching the processing pipeline in [gaddygit]).
For this dataset, we tackle two sub-tasks: Speech Synthesis and Speech Recognition. Speech Synthesis aims at producing audio from EMG data. The full speech production pipeline consists of three steps: a Transduction Model, a Vocoder, and an Automatic Speech Recognition (ASR) model. First, a Transduction model is trained to reconstruct audio target features (Mel-Frequency Cepstral Coefficients, MFCC). It consists of 3 residual blocks (as in [gaddy2021improved]) that downsample the input (1600 samples to 200 samples). The 200 samples are the input to the pre-trained Transformer Encoder; we add relative positional embeddings, following [gaddy2021improved]. A lightweight linear head follows, with an input dimension of $192\times N_{c}$, output dimension of 26 MFCC.
The model is trained using as loss function the mean of pair-wise MFCC distances ($L$) plus an auxiliary phoneme loss ($L_{phon}$)  [gaddy2021improved]. In the case of silent EMG data, DTW is used to align the MFCC features predicted from silent EMG to the sequence of target MFCC features obtained from a parallel vocalized recording. Predicted MFCC audio features are fed to a HiFi-GAN [kong_2020_hifigangenerativeadversarialnetworks] vocoder, which converts the audio features into raw audio waveforms. We adopt the publicly available [gaddygit] pre-trained HiFi-GAN vocoder from Gaddy’s work [gaddy-klein-2020-digital], without any fine-tuning.
For model evaluation, we employ the Wav2Vec2 [baevski_2020_wav2vec20frameworkselfsupervised] ASR model, pretrained on LibriSpeech (English Language) within SpeechBrain [speechbrain].
Speech Recognition consists of recognizing speech (words) without any acoustic output, yielding text as output. Producing text directly from EMG signals is preferable to first synthesizing audio and then applying an ASR system: direct text prediction serves as a more efficient surrogate for human intelligibility, avoids the intermediate audio-generation stage, and yields improved recognition accuracy [gaddy_2022_phdthesis]. Moreover, EMG-to-text models do not require an additional vocoder for speech production, substantially reducing model size and computational load. This makes the approach more suitable for deployment on resource-constrained embedded platforms.
The model architecture is the same as before (Residuals, pre-trained Transformer Encoder and linear layer with softmax). The input to the final linear layers has a dimension of $N_{c}\times 192$, the output dimension is 37 (26 lowercase English letters, 10 digits, and space character). We adopt the Connectionist Temporal Classification (CTC) [graves_2006_connectionist] loss function with a 4-gram language model trained on the LibriSpeech dataset [LibriSpeech]. During inference, beam search is applied to explore the search space and select the paths with the highest likelihood.
The dataset is split as in Gaddy et al. [gaddy2021improved]. Training is performed using both silent and vocalized EMG recordings; testing is performed on silent data. Data are split into training, validation, and test sets using the original index file containing pairs of book and sentence indexes. Each utterance is assigned to one of the three splits. During training, a size-aware sampler is used to group examples, together producing compact batches and a more efficient training stage.
For both tasks, we use the Word Error Rate (WER) as the evaluation metric, defined as
$\text{WER}=\frac{S+D+I}{N},$
where $S$, $D$, and $I$ denote the number of substitution, deletion, and insertion errors, respectively, and $N$ is the total number of words in the reference transcription.
3.2.3 Experimental Settings
All downstream experiments are performed with the same hyperparameters, as per Table 4.
Table 4: Hyperparameters for model fine-tuning.
Hyperparameter Value
Batch size 32
Peak LR $5\times 10^{-4}$
Minimum LR $1\times 10^{-5}$
LR scheduler Cosine
Optimizer $(\beta_{1},\beta_{2})$ AdamW (0.9, 0.98)(0.9,\ 0.98)
Weight decay 0.01
Max training epochs * 50
Warm-up epochs 5
Drop-path 0.10
Layer-wise LR decay 0.90
Label smoothing 0.10
LR: learning rate;
* Early stopping is adopted, with patience of 7 epochs (validation set).
We evaluate the downstream tasks under two different settings: (i) Full-supervised (FS), where the model is trained from scratch on the task-specific dataset without relying on pre-trained weights; (ii) Full Finetuning (FT), where the entire model is fine-tuned using layer-wise learning rate decay to avoid catastrophic forgetting [kenneweg_forgetting_22].
3.3 Model Deployment
We deploy the TinyMyo on the GreenWaves Technologies GAP9 [greenwaves_gap_sdk] processor, a multi-core RISC-V MCU designed for ultra-low-power edge AI. The architecture features a hierarchical memory organization consisting of a large off-chip L3 (HyperRAM), a 1.5 MB on-chip L2 (shared memory), and a 128 kB L1 (Tightly Coupled Data Memory) shared among a cluster of 8 compute cores and a dedicated cluster controller.
To enable the execution of the Transformer-based architecture, which exceeds the on-chip L2 capacity in terms of parameters and intermediate activation maps, we developed a custom deployment toolchain that implements a hierarchical streaming architecture.
Unlike traditional MCU inference engines that require the full model to reside in on-chip memory, our approach treats L2 as a streaming buffer. We implement a multi-level tiling strategy:
  • L3 $\rightarrow$ L2: Large tensors (weights and activations) are stored in external L3 memory. They are retrieved into L2 slabs on demand. For example, in linear and multi-head self-attention layers, weights are streamed slab by slab to prevent memory overflows.
  • L2 $\rightarrow$ L1: Computation is performed by moving data from the L2 slabs into L1 tiles. The cluster controller orchestrates high-bandwidth DMA transfers, while the 8 worker cores execute compute kernels in parallel.
To minimize the overhead of external memory access, we implement a double-buffered pipeline. While the worker cores compute the current slab (N), the cluster controller asynchronously pre-fetches the next slab (N+1) from L3 to L2 or from L2 to L1. This effectively hides the latency of data movement behind the compute-bound operations of the transformer blocks.
We apply INT8 quantization to convert all weights and activations from FP32 to 8-bit fixed-point representations. All intermediate tensors, including those in MHSA (with integer softmax) and GELU/LayerNorm, are stored and processed in INT8 format, with only the LayerNorm scale and shift parameters ($\gamma$ and $\beta$) retained in FP32. The only remaining FP32 operation is the final classification layer, which consumes INT8 inputs and weights but accumulates to FP32 logits to enable a direct comparison with the original FP32 baseline; in principle, this layer could also be executed entirely in INT8.
To maximize the utilization of the limited 1.5 MB L2 memory, we employ an offline liveness analysis pass. Instead of dynamic allocation, the deployment flow calculates a static memory arena. Tensors are assigned offsets within this arena based on their execution lifetime, allowing mutually exclusive buffers (e.g., non-overlapping intermediate feature maps) to share the same physical memory addresses.
To maintain numerical fidelity while avoiding costly floating-point operations, all transformer primitives are implemented in fixed-point arithmetic[kim_2021_bert]. We use integer-only variants of softmax (i-Softmax), layer normalization (i-LayerNorm), and GELU activation (i-GELU), which leverage lookup tables and multiplicative inverses to approximate their floating-point counterparts with minimal error.
As proof of concept, we deploy the Hand Gesture Classification downstream task on the NinaPro DB5 dataset.
Table 5: FLOPs breakdown
Experiment FLOPs
Model Pretraining (Encoder + Decoder) 4.001 G
Ninapro DB5 4.813 G
EPN-612 1.916 G
UCI-EMG 1.916 G
Generic Neuromotor 13.6315 G
Ninapro DB8 4.852 G
Speech Production 1.0956 G
Speech Recognition 1.0921 G
Table 6: Hand Gesture Classification Task — Comparison with State of the Art.
Model Params NinaPro DB5 EPN-612 UCI-EMG
Acc F1 AUROC Acc F1 AUROC Acc F1 AUROC
Moment (as in [chen_2025_physiowave]) 385M 86.41 74.42 N.A. 90.87 90.16 N.A. 90.45 91.75 N.A.
[…]
OTIS (as in [chen_2025_physiowave]) 45M 85.31 72.61 N.A. 87.55 88.03 N.A. 90.62 89.28 N.A.
[…]
PhysioWave-Small [chen_2025_physiowave] 5M 84.78 72.54 N.A. 93.12 93.40 N.A. 90.35 89.51 N.A.
[…]
Notes: PhysioWave [chen_2025_physiowave] use windows of $\approx 1953\text{ms}$. EMG data are upsampled at $2\text{kHz}$ and a fixed window size of 1024 samples is used. WaveFormer [chen_2025_waveformer] uses a fixed window of 1024 samples; no information is available on whether resampling is applied.
FS: model trained from scratch, FT: fine-tuning pretrained model.
4 Results
4.1 Pre-training
Figure 2: Example of reconstruction for a EMG window by the proposed Foundation model. Grey areas correspond to the masked signal regions, white areas to the unmasked.
Figure 2 shows an example of the reconstruction of the EMG signal performed by the model.
With a parameter count of only 3.6 million (with only 0.1% of parameters for the decoder), in line with previous works [chen_2025_waveformer], our model demonstrates that high reconstruction fidelity can be maintained even under a strict parameter budget and low computational requirements A breakdown of the floating-point operations (FLOPs) is available in Table 5.
4.2 Down-stream Tasks Results
1. Hand Gesture Classification
Table 6 reports the results obtained on the Ninapro DB5[ninapro_db5], EPN-612 [benalcazar_emg-epn-612_2020] and the UC Irvine[emg_data_for_gestures_481] datasets, offering a comparison with previous works based on foundation models.
On the NinaPro DB5 dataset, the fine-tuned model trained with a window size of $200\,\text{ms}$ achieves state-of-the-art performance, with an accuracy of $89.41\pm 0.16\,\%$ and an F1-score of $77.97\pm 0.59\,\%$. Fine-tuning the pre-trained model yields an improvement of approximately $7.5\%$ in accuracy compared to training from scratch. Moreover, using a window size of $200\,\text{ms}$ substantially improves performance over $1000\,\text{ms}$ (around $10\%$ for the fine-tuned model). This effect is consistent with the acquisition protocol, which comprises $52$ different movements. Since, during data pre-processing, we did not remove the transition periods between gestures in order to emulate real-life use cases, using shorter windows reduces the likelihood that segments overlapping these transitions are assigned incorrect gesture labels, thereby mitigating label noise. It must be noted that this dataset exhibits a substantial imbalance toward the rest class. To ensure a fair comparison with prior work, we did not apply any rebalancing techniques during training or evaluation. We assume that earlier studies reported the micro accuracy, as no alternative definition was explicitly stated. When computing the macro-averaged accuracy score, the accuracy decreases to $75.80\pm 0.62\,\%$, which reflects the imbalance in the dataset.
On the EPN-612 dataset we achieve SoA results (window size of $1000\text{ms}$, with a global accuracy (FT) of $96.74\pm 0.09\%$, F1-score of $96.74\pm 0.08\%$ and AUROC of $99.70\pm 0.02\%$.
On the UCI-EMG dataset, the best performance is obtained with an input window of $1000\,\text{ms}$, achieving an accuracy of $97.56\pm 0.32$ %, an F1-score of $97.55\pm 0.31$ %, and an AUROC of $99.93\pm 0.02$ %. The performance of the model trained from scratch is comparable to that obtained with the pre-trained TinyMyo encoder, which we attribute to the relative simplicity of this downstream task.
Table 7: Discrete Gesture Classification Task- Comparison with SoA
Model Params CLER Inference Method
LSTM, Kaifosh [generic_neuromotor_interface_2025] (a) 6.4M 0.1819 Full sequence
LSTM, Kaifosh [generic_neuromotor_interface_2025] (b) 6.4M 0.1596 Windowed
[…]
(a): Full sequence results obtained with methods provided by the authors ([generic_neuromotor_interface_2025]) (b): results obtained with the same model as in a, with the windowed-inference method described in the corresponding Methods section. FS: model trained from scratch, FT: fine-tuning pretrained model
For the Generic Neuromotor Interface - discrete gesture task [generic_neuromotor_interface_2025], we achieve a CLER of $0.144\pm 0.004$ on the model trained from scratch, which compares to the 0.1596 (obtained with our windowed inference method, see Sec. 3.2.2) obtained with the LSTM model proposed in the original work, however with $\approx 44\%$ reduction in model size. In this case, the results achieved with the pre-trained Foundation Model are worse ($0.153\pm 0.006$), which we attribute to the intrinsic nature of the problem: to emulate the sequential processing behaviour of the original stacked LSTMs, we imposed a causal mask on the attention mechanism. This constraint restricts our architecture, initially optimized for bidirectional dependency modelling, to a strictly left-to-right context, leading to a degradation in the effectiveness of the pre-trained weights when the future information on which they were originally trained is no longer available. Table 7 reports a summary of the results.
We note that our work is the first to include this dataset among the downstream tasks and to propose an alternative architecture compared to the one employed in the original work.
2. Hand Kinematic Regression
Table 8 reports the result for the Hand Kinematic Regression task.
Table 8: Gesture Regression Task on Ninapro DB8 - Comparison with SoA
Model Params MAE RMSE R2
TEMPONet TCN, Zanghieri [zanghieri_2021_db8] <<500K 6.89
[…]
DNKF, Bao [BAO202188]∗∗ 13.5
[…]
5 DoFs, subject-specific model; ∗∗ 3 DoFs, subject-specific model;
∗∗∗ 5 DoFss, across-subject model;
FS: model trained from scratch, FT: fine-tuning pretrained model
We obtain a MAE of $8.77\pm 0.12^{\circ}$, with a window size of $1000\text{ms}$ . This value is higher than the result reported by Zanghieri et al. [zanghieri_2021_db8] (MAE $6.89^{\circ}$), primarily due to differences in both the training and evaluation protocols. In our setting, the model is trained in an across-subject regime using data aggregated from all participants, while prior work trained subject-specific models, i.e., one model per individual. Learning a single model that generalizes across subjects is substantially more challenging because of inter-subject variability in muscle physiology, electrode placement, and EMG signal morphology.
3. Speech Recognition
Table 9: Silent Speech Synthesis Task – Comparison with SoA
Transduction Model Vocoder ASR WER
Method Model Params [M] Model Params [M]
Transformer, Gaddy [gaddy2021improved] CNN + Transformer 54 HiFi-GAN 13.92 DeepSpeech 36%
Diff-ETS, Ren [ren_speech_24] as [gaddy2021improved] + Diffusion Probabilistic Model >>54 (a) HiFi-GAN 13.92 Whisper Medium (EN) 32%
SU-E2S, Scheck. [Scheck2023SUE2S] HuBERT-Soft N.A. HiFi-GAN 13.92 Whisper Medium (EN) 40%
DiffMV-ETS [Scheck2025DiffMVETSDM] DiffMV-ETS $\approx$ 66 (b) BigVGAN 112 Distil-Whisper-v3 N.A.
EMGVox-GAN [SUALIHEEN2025101754] as [gaddy2021improved] as [gaddy2021improved] EMGVox-GAN 12 N.A. 36%
SWS, Lee [lee_speech_2025] as [gaddy2021improved] 54 Sovits-SVC N.A. Whisper Medium (EN) 38%
[…]
(a) = estimated. The authors employ the same Transduction model as in Gaddy [gaddy2021improved], followed by a Diffusion Model (including residual blocks, attention and convolutional layers)
(b) = estimated as sum of speaker encoder (6M), prior encoder (30M), and diffusion decoder (30M)
FS: model trained from scratch, FT: fine-tuning pretrained model
Table 9 reports the results for the Speech Synthesis Task.
Our model achieves a WER (FT) of $33.54\pm 1.12\%$, in line with previous works (SoA WER is 32%, achieved by Ren et al. [ren_speech_24]). Compared to previous solutions, our Transduction model offers a significant reduction in parameter size, with 62.5% reduction compared to Scheck et al. [Scheck2023SUE2S] and 91.7% reduction compared to Gaddy et al. [gaddy2021improved].
The reported number of parameters accounts only for Transduction model (encoder and model head). The number of parameters for the pre-trained Hi-Fi Gan vocoder is 13.92M (same as in [gaddy2021improved, ren_speech_24]).
Table 10: Silent Speech Recognition Task - Comparison with SoA
Method Params WER
Transformer, Gaddy [gaddy_2022_phdthesis] 54 M 28.8%
MONA, Benster [benster_2024_mona] (a) n.a. 22.2%
MONA LISA, [benster_2024_mona] (a) n.a. 12.2%
[…]
(a): Multimodal model
FT: fine-tuning pretrained model
Table 10 reports the results for the Speech Recognition Task. Given the superior performance of the fine-tuned model compared to training from scratch in the Speech Production task, and the identical architecture adopted for the vocoder, in this case, we run the experiments only for the full fine-tuning setting. We achieve a WER of $33.95\%\pm 0.97\%$, higher compared to 28.8% reported of Gaddy et al. [gaddy_2022_phdthesis], however, with a 91.7% reduction in parameter cont.
The MONA LISA model [benster_2024_mona] achieves a WER of 12.2%. However, this performance relies on a multimodal training setup, where both EMG and audio data are jointly used. Moreover, MONA LISA is a large-scale language model, which poses practical limitations for real-time deployment and edge applications due to its substantial computational footprint and memory requirements.
4.3 Model Deployment
We deploy TinyMyo on the GAP9 microcontroller, as described in Section 3.3. The model comprises 3.56 million parameters, with the majority residing in the 8 transformer blocks. Each block contains a multi-head self-attention (MHSA) layer with 3 heads and a feed-forward network (FFN) with an expansion ratio of 4.
Table 11 reports a computational cost breakdown of the deployed model for one Transformer Block. We can observe that the computational cost is dominated by the MHSA QK/AV operations (40% of total MACs) and the MLP projections (40% of total MACs). In particular, the QK and AV terms scale as $\mathcal{O}(N^{2}d)$ with the sequence length $N=800$, making them the asymptotic computational bottleneck of the Transformer block.
With GAP9 operating at 370 MHz, utilizing all 9 cores in the cluster to execute TinyMyo computations in parallel, we achieve an inference in around $12.2\text{seconds}$ at only $0.44\text{J}$ with an average power-envelope of $36.45\text{mW}$.
Table 11: Computational Cost per Transformer Block
Component MACs % of Total
MHSA Q/K/V projections 88M 15%
MHSA QK scores 123M 20%
MHSA AV context 123M 20%
MHSA output projection 29M 5%
MLP FC1 118M 20%
MLP FC2 118M 20%
Notes: QK scores and AV context is (800 $\times$ 800 $\times$ 3 heads). FC1 and FC2 are $192\to 768$ and $768\to 192$ with $4\times$ MLP expansion factor.
5 Conclusion
In this work, we presented TinyMyo, the first compact EMG FM capable of broad generalization across diverse downstream tasks and real-time edge deployment on an ultra-low-power microcontroller.
TinyMyo is based on a Transformer-encoder architecture, with a size of 3.6 million parameters. The model is pre-trained using heterogeneous EMG data, collected from platforms with different numbers of channels and from diverse body locations.
TinyMyo achieves SoA performance across multiple downstream tasks. For gesture recognition, we obtain SoA accuracy on the NinaPro DB5 ($89.41\pm 0.16\,\%$), EPN-612 ($96.74\pm 0.09\,\%$), and UCI-EMG ($97.56\pm 0.32\,\%$) datasets. Furthermore, on the Generic Neuromotor Interface – discrete-gesture dataset, TinyMyo achieves a CLER of $0.153\pm 0.006$ with an approximate $44\,\%$ reduction in model size compared to the original work.
On the hand kinematic regression task, we achieve a MAE of $8.77\pm 0.12$ with a subject-agnostic model. On the speech production task, we obtain a WER of $33.54\pm 1.12\,\%$, which is comparable to the current SoA, while reducing the model size by approximately $91\,\%$. Finally, we demonstrate that TinyMyo can also adapt to the speech recognition task, achieving a WER of $33.95\pm 0.97\%$.
TinyMyo is also the first FM for EMG deployed on the edge, achieving an inference in $12.2\text{s}$ at $0.44\text{J}$, with an average power-envelope of $36.45\text{mW}$.
We identify the following limitations, which will be addressed as part of future work:
  • The datasets used for pre-training consist primarily of EMG recordings from the forearm and wrist during gesture-related tasks. Data from additional body locations (e.g., lower-limb muscle groups) should be included to improve anatomical generalization.
  • Both the pre-training and downstream datasets were collected in controlled laboratory environments. Additional research is needed to evaluate the robustness of the proposed FM under realistic conditions, such as data corrupted by motion artifacts or environmental noise.
  • Further studies are required to assess the model’s generalization capabilities in zero-shot settings and to quantify the amount of labeled data needed to achieve strong performance on novel downstream tasks.
  • Despite demonstrating, to the best of our knowledge, the first deployment of an FM for EMG processing on an ultra–low-power microcontroller, the inference time remains too high for seamless use in real-world applications. Future work should therefore focus on optimizing the network architecture to minimize the computation needed for inference. This could be achieved by doing an ablation study on reducing the sequence length, by adopting a cheaper attention mechanism [dimofte_2025_cerebro] or exploring more aggressive quantization techniques (e.g.: sub 8-bit [dong_2023_packqvit]. More efficient architectures, such as state-space models (e.g., MAMBA [gu_2024_mamba], FEMBA [tegon_2025_femba]), whose computational cost scales linearly with the context length, would also be worth exploring. Moreover, distributing the computation over multiple GAP9 nodes or deploying it on more powerful platforms such as Siracusa [prasad_2024_siracusa] is expected to yield substantial speed-ups.
Architectural improvements and refinements in the training pipeline also represent an important future direction. We aim to explore alternative pooling mechanisms (e.g., attention-based) to weight per-channel contribution more precisely.
By open-sourcing111after review our pre-trained models and the architectures used for downstream evaluations, we aim to provide a versatile resource that can accelerate future research and serve as a foundation for the broader EMG community. Our intention is that other researchers will extend this backbone with additional datasets, improved training strategies, and architectural variations, aligning with the community-wide effort recently advocated by Eddy et al. [eddy_2025_raising].
Acknowledgment
The authors thank Chen Yanlong (ETH Zürich) for technical support and Lorenzo Lamberti (ETH Zürich) for fruitful discussions.

### v2

@@
 Abstract
-Surface electromyography (EMG) is a non-invasive sensing modality used in several domains, including biomechanics, rehabilitation, prosthetic control, and emerging human-machine interaction paradigms.
-Despite decades of use, significant challenges remain in achieving robust generalization across subjects, recording systems, and acquisition protocols.
-To tackle these challenges, foundation models (FMs) are gaining traction when targeting end-to-end applications based on EMG signals.
-Yet, existing EMG FMs remain limited to single downstream tasks and lack deployability on embedded platforms.
-In this work, we present TinyMyo, a lightweight FM based on a Transformer encoder architecture.
-The model is pre-trained in a self-supervised manner on publicly available datasets and achieves high reconstruction fidelity with only 3.6M parameters.
-With minimal task-specific head adaptations, the same backbone is used to tackle multiple downstream tasks, leveraging datasets acquired from diverse sensing locations and hardware platforms.
-We demonstrate generalization across hand gesture classification, hand kinematic regression, speech production and recognition, with performance comparable to or surpassing the state of the art (SoA), and model size below 5M parameters.
-We achieve SoA results compared to previous FM-based works on the NinaPro DB5 ($89.4\pm 0.16\%$), UCI-EMG ($97.56\pm 0.32\%$), and EPN-612 ($96.74\pm 0.09\%$) datasets.
-We report, to the best of our knowledge, the first deployment of an EMG FM on an ultra-low-power microcontroller (GAP9), achieving an average power envelope of $36.45\text{mW}$.
-By open-sourcing the pre-trained and the downstream task architectures (https://github.com/pulp-bio/BioFoundation), we aim to provide a flexible resource that can accelerate future research and serve as a common foundation for the EMG community.
+Objective: Surface electromyography (EMG) is a non-invasive sensing modality widely used in biomechanics, rehabilitation, prosthetic control, and human-machine interfaces.
+Despite decades of use, achieving robust generalization across subjects, recording systems, and acquisition protocols remains challenging.
+While foundation models (FMs) are gaining traction for EMG, existing approaches remain limited to single downstream tasks and lack deployability on embedded platforms.
+This work addresses these limitations.
+Methods: We present TinyMyo, a lightweight FM based on a Transformer encoder architecture.
+The model is pre-trained in a self-supervised manner using masked reconstruction on publicly available datasets.
+With only 3.6M parameters, TinyMyo is designed to support multiple downstream tasks through minimal task-specific head adaptations.
+Results: We demonstrate generalization across hand gesture classification, hand kinematic regression, speech production and speech recognition, with performance comparable to or surpassing the state of the art (SoA), and model size below 5M parameters.
+We achieve SoA results compared to previous FM-based works on the NinaPro DB5 (89.4%), UCI-EMG (97.56%), and EPN-612 (96.74%) datasets.
+We demostrate the first-time deployment of an EMG FM on an ultra-low power microcontroller (GAP9), with an inference time of 0.785 s, energy of 44.91 mJ and power envelope of 57.18 mW.
+Conclusion: TinyMyo demonstrates that compact, self-supervised EMG FM can guarantee strong generalization across multiple downstream tasks while remaining compatible with low-power edge devices.
+Significance: TinyMyo is the first EMG FM for ultra-low-power edge-devices, enabling scalable and energy-efficient sensing for motor intent decoding, neuromuscular assessment, and biosignal-driven human–machine interaction.
 {IEEEkeywords}
@@
 1 Introduction
-Electromyography (EMG) measures variations in the electrical activity of muscles during contractions, reflecting the motor-neuron activation patterns [farina_2023_rev].
-EMG can be acquired using invasive methods (needle EMG), suitable for a detailed characterization of individual motor units, or using non-invasive, surface-based techniques, to capture the global activity of muscle fibres  [farina_2023_rev, avila_2023_rev].
-Surface EMG has found extensive application across a wide range of domains, including biomechanical analysis and rehabilitation [avila_2023_rev], human–robot interaction [xiong_emgrobot_2024], prosthetic control [cimolato_2022emglower, marinelli_2023emgupper], and emerging human–computer interaction interfaces [generic_neuromotor_interface_2025].
+Electromyography (EMG) measures variations in the electrical activity of muscles during contractions, reflecting the motor-neuron activation patterns [15].
+EMG can be acquired using invasive methods (needle EMG), suitable for a detailed characterization of individual motor units, or using non-invasive, surface-based techniques, to capture the global activity of muscle fibres  [15, 3].
+Surface EMG has found extensive application across a wide range of domains, including biomechanical analysis and rehabilitation [3], human–robot interaction [55], prosthetic control [35], and emerging human–computer interaction interfaces [26].
 However, despite decades of widespread use, many important challenges remain.
-First, EMG signals are highly affected by noise, motion artifacts, and cross-talk from adjacent muscles [boyer_emg_noise].
-Also, acquisition systems present large variability in terms of number and arrangement of electrodes, sensor geometry, sampling rates, amplification stages, and filtering pipelines [merletti_2020_tutorial].
-These sources of variability pose challenges to the design of models that can generalize across heterogeneous hardware configurations.
-In addition, EMG recordings are strongly influenced by both inter-subject and intra-subject variability [xiong_2021_review].
-Differences in anatomy, subcutaneous fat, electrode placement, skin impedance, and day-to-day physiological or behavioural factors lead to substantial shifts in the EMG distribution, posing an additional challenge to generalization across users or recording sessions [ni_2024_survey].
+EMG signals are strongly affected by noise, motion artifacts, cross-talk, and acquisition variability across electrode layouts, hardware platforms, and acquisition pipelines [8, 36].
+In addition, both inter-subject and intra-subject factors [56], including anatomy, electrode placement, skin impedance, and physiological variability lead to substantial distribution shifts, hindering generalization [37].
 Foundation Models (FM) offer a great promise to address these challenges.
-Inspired by the success of self-supervised pre-training in vision, language, and speech, as well as by emerging applications in biosignals [gu_2025_foundationsurvery], the core hypothesis is that large-scale datasets can be leveraged to learn high-level, platform and subject-agnostic signal representations.
+Inspired by the success of self-supervised pre-training in vision, language, and speech, as well as by emerging applications in biosignals [23], the core hypothesis is that large-scale datasets can be leveraged to learn high-level, platform and subject-agnostic signal representations.
 For EMG, such models could provide a unified backbone for many downstream tasks, building upon a robust, device-agnostic feature extraction, improving generalization across devices and subjects.
-However, progress in FMs of EMG has been limited by the scarcity of large sEMG corpora, and only with recent large-scale datasets (collected from thousands of users) [salter_emg2pose_2024] the exploration of generalized FMs of EMG became viable.
+However, progress in FMs of EMG has been limited by the scarcity of large sEMG corpora, and only with recent large-scale datasets (collected from thousands of users) [43] the exploration of generalized FMs of EMG became viable.
 In this context, existing EMG FMs present significant limitations.
-Current models mostly focus on a single task (typically gesture classification) [chen_2025_physiowave, chen_2025_waveformer].
+Current models mostly focus on a single task (typically gesture classification) [10, 11].
 Furthermore, the size of EMG FMs remains large, and no prior work demonstrated the deployment of an EMG FM on embedded hardware.
-This motivates the need for efficient FM architectures that meet the memory and computational constraints of embedded devices [benatti_2025_emgreview].
-We address these limitations by proposing the following contribution.
+This motivates the need for efficient FM architectures that meet the memory and computational constraints of embedded devices.
+We address these limitations via the following contribution.
   • We present TinyMyo, a lightweight Transformer-based encoder operating directly on EMG signals in the time domain.
@@
   • Unlike previous EMG FMs, which are typically designed around a single downstream task or rely on large, domain-agnostic encoders, TinyMyo generalizes across heterogeneous sEMG datasets and sensor configurations.
-Our architecture follows the general Transformer design adopted in prior EMG works [chen_2025_physiowave, chen_2025_waveformer], formulated in a unified EMG FM able to support a wide range of downstream tasks using the same pretrained backbone.
-We demonstrate the versatility of the pre-trained encoder through minimal task-specific head adaptations across diverse downstream tasks, including hand-gesture classification on the Ninapro DB5 [ninapro_db5], EPN-612 [benalcazar_emg-epn-612_2020], UC Irvine [emg_data_for_gestures_481] and Meta’s Generic Neuromotor Interface [generic_neuromotor_interface_2025] datasets, hand-kinematic regression on the Ninapro DB8 [ninapro_db8], and speech production and discrimination on the Gaddy Silent Speech Dataset [gaddy-klein-2020-digital].
+Our architecture follows the general Transformer design adopted in prior EMG works [10, 11], formulated in a unified EMG FM able to support a wide range of downstream tasks using the same pretrained backbone.
+We demonstrate the versatility of the pre-trained encoder through minimal task-specific head adaptations across diverse downstream tasks, including hand-gesture classification on the Ninapro DB5 [39], EPN-612 [6], UC Irvine [32] and Meta’s Generic Neuromotor Interface [26] datasets, hand-kinematic regression on the Ninapro DB8 [31], and speech production and discrimination on the Gaddy Silent Speech Dataset [16].
 Our model achieves competitive performances across all downstream tasks, either surpassing SoA or guaranteeing a comparable accuracy with a sub-5M class model size.
-  • We demonstrate for the first time the deployment of an EMG FM on an ultra-low-power microcontroller (GAP9), achieving an average power envelope of $36.45\text{mW}$.
-  • We open source the model and weights to enable reproducibility and further research by the community.
+  • We demonstrate for the first time the deployment of an EMG FM on an ultra-low-power microcontroller (GAP9), achieving an inference time of $0.785~\text{s}$, energy of $44.91~\text{mJ}$ and power envelope of $57.18~\text{mW}$.
+  • We open source the model and the weights to enable reproducibility and further research by the community (https://github.com/pulp-bio/BioFoundation).
 2 Related Works
-Several efforts have been made to develop FMs for biosignals, and most of the works focused on electrocardiography (ECG), photoplethysmography (PPG), and electroencephalography (EEG) [ni_2024_survey].
+Several efforts have been made to develop FMs for biosignals, and most of the works focused on electrocardiography (ECG), photoplethysmography (PPG), and electroencephalography (EEG) [37].
 In contrast, research dedicated to EMG FMs remains limited.
-Turgut et al.
-[turgut_2024_otis] introduced OTiS, a foundation model with 45 M parameters pre-trained on a broad collection of modalities, including ECG, temperature, audio, electromechanical signals, EEG, and financial time series.
-Their work demonstrated the feasibility of handling heterogeneous time-series data, including diverse input types, sampling rates, and channel configurations, within a unified pre-training framework.
+OTiS [50] is a 45 M parameters FM pre-trained on a broad collection of modalities, which demonstrated the feasibility of handling heterogeneous time-series data, including diverse input types, sampling rates, and channel configurations, within a unified pre-training framework.
 In the original study, EMG is used exclusively for multi-class muscular disease classification.
 Subsequent works by Cheng et al.
-[chen_2025_physiowave, chen_2025_waveformer] evaluated the same model on hand-gesture classification using the EPN612 [benalcazar_emg-epn-612_2020], NinaPro DB5 [ninapro_db5], NinaPro DB6 [ninapro_db6], and UCI EMG [emg_data_for_gestures_481] datasets, demonstrating high accuracy across all datasets.
-Goswami et al.
-proposed MOMENT [goswami_2024_moment], a 350-M parameter FM pre-trained on a wide range of domains and temporal resolutions, including electrical signals, traffic, weather, disease progression, motion capture, simulated data, and ECG (among biopotentials).
-In the original work, EMG is not included in the pre-training corpus nor in the downstream evaluations.
-However, recent studies by Cheng et al.
-[chen_2025_waveformer, chen_2025_physiowave] assessed MOMENT on the hand-gesture classification task using the EPN612, NinaPro DB5, NinaPro DB6, and UCI EMG datasets, demonstrating the applicability of MOMENT as an EMG FM.
-PhysioWave [chen_2025_physiowave] introduced a modality-specific FM for EMG and ECG.
-The authors further propose a unified multimodal framework combining these encoders with existing pre-trained EEG models, leveraging a multi-scale wavelet-transformer architecture, using wavelet decomposition to capture time-frequency features at different scales.
-For the EMG-specific model, the primary downstream task is hand-gesture classification, evaluated on NinaPro DB5, EPN612, and UCI EMG.
+[10, 11] evaluated the same model on hand-gesture classification using the EPN612 [6], NinaPro DB5 [39], NinaPro DB6 [38], and UCI EMG [32] datasets, demonstrating high accuracy across all datasets.
+MOMENT [19] is a 350-M parameter FM pre-trained on a wide range of domains and temporal resolutions.
+While the original work did not consider EMG, recent works [11, 10] evaluated it for hand-gesture classification on the EPN612, NinaPro DB5, NinaPro DB6, and UCI EMG datasets.
+PhysioWave [10] introduced a modality-specific FM for EMG, primarily evaluated on hand-gesture classification using the NinaPro DB5, EPN612, and UCI EMG.
 PhysioWave achieved a performance comparable to Moment and OTIS with a much smaller parameter count (5M for the smaller model).
-The same authors proposed WaveFormer [chen_2025_waveformer], a lightweight Transformer model tailored for sEMG-based gesture recognition, with a model size of only 3.1 M parameters.
-To ensure the extraction of frequency-domain features remains compact, the model integrates a learnable wavelet transform based on multi-level wavelet decomposition and depthwise separable convolutions.
-Downstream evaluations are conducted on the EPN612, NinaPro DB5, NinaPro DB6, and UCI EMG datasets, further improving the performance over PhysioWave.
-More recently, Jiang et al.
-[jiang_2025_towards] introduced PhysioOmni, a multi-modal FM designed to handle a wide range of physiological signals such as EEG, ECG, EOG and EMG.
-For each modality, the authors use the same encoder as in LaBraM [jiang_2024_labram] (5.85.8M parameters per modality, considering its smallest-Base version).
+WaveFormer [11] further refined this approach with a lightweight Transformer model (3.1 M parameters) tailored for sEMG-based gesture recognition.
+WaveFormer improved performance over PhysioWave on the EPN612, NinaPro DB5, NinaPro DB6, and UCI EMG datasets.
+Jiang et al.
+[24] introduced PhysioOmni, a multi-modal FM covering EEG, ECG, EOG and EMG, based on the same encoder as in LaBraM [25] (5.85.8M parameters per modality, considering its smallest-Base version).
 They adopt a decoupled multimodal tokenizer and masked signal modelling to ensure robustness under arbitrary missing modalities during inference.
 During pre-training, PhysioOmni relies on a fixed set of modalities (EEG, EOG, and EMG), which makes the architecture unsuitable for pre-training when one modality is missing.
-For downstream evaluation, the authors include 5-class sleep stage classification on the HMC [alvarez_2021_hmc] dataset and gait kinematic regression on the FBM [brantley_2018_fbm] dataset.
-In this context, existing EMG FMs appear limited by the small number of explored downstream tasks, as well as by the lack of deployability on embedded platforms.
+Downstream evaluations include sleep stage classification [2] and gait kinematic regression[9].
+In this context, existing EMG FMs are limited by the small number of explored downstream tasks, as well as by the lack of deployability on embedded platforms.
 TinyMyo is the first EMG FM to cover a broader range of downstream tasks with SoA performance, as will be extensively described in Section 4, while also being deployed on an ultra-low-power MCU.
@@
 The output is text for the Speech Recognition task.
-For the the Speech Production task, a vocoder (HiFi-GAN) is added to produce audio from the predicted MFCC features.
+For the Speech Production task, a vocoder (HiFi-GAN) is added to produce audio from the predicted MFCC features.
 (c) Model Deployment: the architecture for the hand-gesture classification task is implemented on the GAP9 microcontroller.
@@
 3.1.1 Tokenization
-TinyMyo adopts a Channel-Independent patching strategy, following methodologies effectively applied in time-series forecasting [nietime] and vision [bao2024_channel_visiontransformersimage].
-Unlike other recent biosignal (EEG) foundation models that mix spatial information during the tokenization phase via 2D convolutions that span multiple channels (e.g., FEMBA uses a $2\times 16$ kernel [tegon_2025_femba]) or project them into a unified latent space (e.g., LUNA [doner2025luna]), we strictly maintain per-channel granularity.
+We adopt a channel-independent patching strategy, preserving per-channel granularity at the embedding stage.
+Unlike EEG models that mix spatial information during tokenization via 2D convolutions spanning multiple channels [49] or project them into a unified latent space [13], we embed channels independently using a $1\times L$ kernel, and the subsequent transformer layers learn inter-channel dependencies via attention [12].
 This design choice is grounded in the physiological nature of EMG recordings, where each channel corresponds to a distinct anatomical location.
-By treating the channels independently at the embedding stage (using a $1\times L$ kernel), the model allows the subsequent transformer layers to learn inter-channel dependencies via attention, a strategy proven effective for biosignal decoding [dimofte_2025_cerebro].
-Formally, let the input EMG waveform be denoted as $\mathbf{X}\in\mathbb{R}^{T\times C}$, with $T$ timesteps and $C$ channels.
-We apply a sliding window with patch length $L$ and stride $S$ independently to each channel.
+Formally, given an input EMG signal $\mathbf{X}\in\mathbb{R}^{T\times C}$, with $T$ timesteps and $C$ channels, we apply a sliding window with patch length $L$ and stride $S$ independently to each channel.
 This yields a grid of patches $\mathbf{P}\in\mathbb{R}^{C\times N_{p}\times L}$, where $N_{p}=\lfloor(T-L)/S\rfloor+1$ is the number of temporal patches per channel.
 We employ a window of $T=1000$ timesteps, $L=S=20$, and $C=16$, yielding to $50$ tokens per channel ($N_{p}$) and a total sequence length of $800$ tokens ($N$).
-To form the input sequence for the Transformer, we flatten the channel and temporal dimensions, resulting in a total sequence length of $N=C\times N_{p}$.
-Each patch $\mathbf{P}_{c,i}$ (where $c$ is the channel index and $i$ is the patch index) is mapped to a latent embedding of dimension $d_{e}$ via a shared learnable linear projection $\mathbf{W}_{\mathrm{proj}}\in\mathbb{R}^{d_{e}\times L}$.
-The resulting token sequence $\mathbf{E}\in\mathbb{R}^{N\times d_{e}}$ consists of embeddings:
-$\mathbf{E}_{k}=\mathbf{W}_{\mathrm{proj}}\,\mathbf{P}_{c,i}^{\top},$ (1)
-where the token index $k$ corresponds to the flattened arrangement of channel-time pairs $(c,i)$.
-We do not add learnable positional embeddings; instead, we inject positional information via Rotary Position Embeddings (RoPE) [su_2023_roformerenhancedtransformerrotary] within the attention blocks.
+As input sequence for the Transformer, we flatten the channel and temporal dimensions, resulting in a total sequence length of $N=C\times N_{p}$.
+Each patch $\mathbf{P}_{c,i}$ ($c$ is the channel index and $i$ is the patch index) is mapped to a latent embedding of dimension $d_{e}$ via a shared learnable linear projection $\mathbf{W}_{\mathrm{proj}}\in\mathbb{R}^{d_{e}\times L}$.
+The resulting token sequence $\mathbf{E}\in\mathbb{R}^{N\times d_{e}}$ consists of embeddings $\mathbf{E}_{k}=\mathbf{W}_{\mathrm{proj}}\,\mathbf{P}_{c,i}^{\top}$ where the token index $k$ corresponds to the flattened arrangement of channel-time pairs $(c,i)$.
+We do not add learnable positional embeddings; instead, we inject positional information via Rotary Position Embeddings (RoPE) [47] within the attention blocks.
 3.1.2 Architecture
@@
 The Encoder (see Fig.
-1) consists of 8 identical layers, based on pre-LayerNorm Transformer encoder [vaswani_2023_attentionneed, xiong_2020_layernormalizationtransformerarchitecture], with 3 attention heads and an embedding dimension of 192, analogous to the configuration used in EfficientViT [liu2023efficientvitmemoryefficientvision].
-We adopt a RoPE approach, which introduces position information by rotating query and key vectors in 2D planes instead of adding a position vector; such a method has already been adopted for biosignals [wang_2024_eegpt, chen_2025_physiowave, chen_2025_waveformer].
+1) comprises 8 layers with an embedding dimension of 192 and 3 attention heads.
+We utilize an MLP ratio of 4.0 (size 768), QKV bias (without QK norm), and a dropout rate of 0.1 for attention, projection, and drop-path.
+The lightweight Decoder shares the embedding dimension of 192.
+We adopt a RoPE approach, which introduces position information by rotating query and key vectors in 2D planes instead of adding a position vector; such a method has already been adopted for biosignals [52, 10, 11].
 During pre-training, the Encoder’s output is passed to a lightweight Decoder (3.9k parameters), consisting of a linear layer that outputs a sequence $\mathbf{\hat{P}}$, which is a reconstruction of the original patch sequence $\mathbf{P}$.
-Our approach draws inspiration from SimMIM [xie2022simmimsimpleframeworkmasked], which demonstrates that a minimal reconstruction head (e.g., a single linear layer) is sufficient for effective self-supervised pre-training.
+Our approach draws inspiration from SimMIM [54], which demonstrates that a minimal reconstruction head (e.g., a single linear layer) is sufficient for effective self-supervised pre-training.
 The use of a lightweight decoder is intentional: placing the full reconstruction burden on the encoder forces it to learn semantically rich and task-agnostic representations, rather than relying on a high-capacity decoder to compensate for missing information.
-This design strategy is supported by recent survey analyses in the masked image modeling (MIM) literature: shallow decoders tend to limit shortcut learning and encourage the formation of stronger latent representations within the encoder [zhang2022surveymaskedautoencoderselfsupervised].
+This design strategy is supported by recent survey analyses in the masked image modeling literature: shallow decoders limit shortcut learning and encourage the formation of stronger latent representations within the encoder [59].
 As a result, the learned features exhibit improved generalization across downstream tasks and recording conditions.
-Further details on the model’s parameters are reported in Table 1 (left).
 3.1.3 Pre-training objectives
-We adopt a random-masking approach [xie2022simmimsimpleframeworkmasked], where a random subset $\mathcal{M}$ of tokens is replaced by a learnable $\mathcal{[MASK]}$ token.
+We adopt a random-masking approach [54], where a random subset $\mathcal{M}$ of tokens is replaced by a learnable $\mathcal{[MASK]}$ token.
 Applying the masking over each patch enforces contextual inference over tens of milliseconds (physiologically meaningful burst segments), rather than trivial gap filling.
-We adopt a masking ratio of $50\%$ as in previous works [dimofte_2025_cerebro, tegon_2025_femba].
-We employ the Smooth L1 Loss function as reconstruction objective between masked and original patches, as in [chen_2025_physiowave]. We define the following loss components:
+We adopt a masking ratio of $50\%$ as in previous works [12, 49].
+We employ the Smooth L1 Loss function as reconstruction objective between masked and original patches, as in [10]. We define the following loss components:
 $\mathcal{L}_{\mathrm{masked}}=\frac{1}{|\mathcal{MASK}|}\sum_{(c,i)\in\mathcal{MASK}}\mathrm{SmoothL1}(\mathbf{P}_{c,i},\mathbf{\hat{P}}_{c,i}),$ (1)
@@
 $\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{masked}}+0.1\cdot\mathcal{L}_{\mathrm{visible}}$ (3)
-Additional details on pre-training set-up are provided in Table 1 (right).
-Table 1: Hyperparameters for masked EMG pre-training.
-Hyperparameter Value
-Transformer Encoder
-Timesteps 1000
-Patch size $(H{\times}W)$ $1\times 20$
-Channels 16
-Embed dim 192
-Layers 8
-Heads 3
-QKV bias Yes
-QK norm No
-MLP ratio 4.0
-MLP size 768
-Attn drop 0.1
-Proj drop 0.1
-Drop-path 0.1
-Decoder
-Decoder embed dim 192
-Setup Value
-Batch size 512
-Peak LR $1\!\times\!10^{-4}$
-Minimum LR $1\!\times\!10^{-6}$
-Optimizer $(\beta_{1},\beta_{2})$ AdamW (0.9,0.98)
-LR schedule Cosine
-Weight decay 0.01
-Epochs / warm-up 50 / 10
-Grad accum batches 8
-Grad clip 3
-Mask ratio 0.5
-Max seq length 1000
-LR: learning rate
 All experiments are implemented in Python 3.10 using PyTorch Lightning and Hydra for modularity, configurability, and reproducibility.
 The proposed foundation model is trained on the CSCS Alps HPC infrastructure using NVIDIA GH200 GPUs, employing a single node with 4x NVIDIA GH200 GPUs in Distributed Data Parallel (DDP) mode.
+We train using the AdamW optimizer ($\beta_{1}=0.9,\beta_{2}=0.98$, weight decay 0.01) with a global batch size of 512, utilizing gradient accumulation over 8 batches and a clipping threshold of 3.
+The learning rate follows a cosine schedule for 50 epochs (10 warm-up epochs), peaking at $1\times 10^{-4}$ (minimum $1\times 10^{-6}$).
+The model processes sequences of 1000 timesteps with a masking ratio of 0.50 .
 3.1.4 Pre-training Datasets
-For model pre-training, we employ three publicly available datasets, with a summary provided in Table 2.
-Table 2: Pre-training EMG datasets.
-Dataset #Sub $f_{s}\text{[Hz]}$ #Ch Size [GB] Placement
-Ninapro DB6 [ninapro_db6] 10 2000 14 20.3 Forearm
-Ninapro DB7 [ninapro_db6] 22 2000 12 30.9 Forearm
-EMG2Pose [salter_emg2pose_2024] 192 2000 16 431 Wrist
-Ninapro DB6 [ninapro_db6] contains sEMG from 10 subjects performing 7 hand grasp types, each repeated 12 times over 5 separate recording days.
+Ninapro DB6 [38] contains sEMG from 10 subjects performing 7 hand grasp types, each repeated 12 times over 5 separate recording days.
 Signals were acquired with Trigno sensors (14 channels, $2~\text{KHz}$ sampling rate), placed on the forearm.
 The total dataset size is $20.3~\text{GB}$.
-Ninapro DB7[ninapro_db7] contains data from 20 intact subjects and 2 transradial amputees, performing 40 movements spanning isolated finger/wrist actions and grasp patterns.
+Ninapro DB7[30] contains data from 20 intact subjects and 2 transradial amputees, performing 40 movements spanning isolated finger/wrist actions and grasp patterns.
 Data were collected using Trigno sensors (12 channels, $2~\text{kHz}$ sampling rate), placed on the forearm.
 The total dataset size is $30.9~\text{GB}$.
-EMG2Pose[salter_emg2pose_2024] contains data from 193 participants, performing 29 different hand motions.
+EMG2Pose[43] contains data from 193 participants, performing 29 different hand motions.
 Data were recorded with a custom wristband (16 channels, $2~\text{kHz}$ sampling rate).
@@
 3.1.5 Pre-training Data Processing
-EMG data are pre-processed using a Butterworth band-pass filter ($4^{th}$ order, $20-450~\text{Hz}$), followed by a notch filter ($50~\text{Hz}$), as in [chen_2025_physiowave].
-Min-Max, channel-wise normalization ([1,1][-1,1] range) is applied.
-Each recording is then segmented into fixed-length windows of 1000 samples ($500~\text{ms}$ at $2~\text{kHz}$) with 50% overlap, providing sufficient temporal context [smith_emgwindow_2011][delafuente_emgwindow_2021] while controlling computational load.
+EMG data are band-pass filtered ($4^{th}$ order Butterworth, $20-450~\text{Hz}$), notch filtered ($50~\text{Hz}$), normalized to [1,1][-1,1], and segmented into 1000-sample windows ($500~\text{ms}$ at $2~\text{kHz}$) with 50% overlap, providing sufficient temporal context [46] while controlling computational load.
 Datasets with fewer than 16 channels are zero-padded. Given that each training batch might contain data from each dataset, this approach ensures consistency in the number of channels per batch.
-We did not apply any data augmentation in line with previous works [chen_2025_physiowave, chen_2025_waveformer] which already showed SoA performances.
-Additionally, given the small model size (3.6 M parameters) and the limited effect of data augmentation shown in scaling-laws regimes [kaplan2020scalinglawsneurallanguage], we rely on the sample-efficiency of the architecture rather than increasing artificially the data variance.
+We did not apply any data augmentation due to the limited effect in scaling-laws regimes [27] and in line with previous works [10, 11].
 3.2 Downstream Tasks
 3.2.1 Pipeline Overview
-For all downstream tasks, the Encoder architecture described in Sec. 3.1.2 is retained without modification, while the Decoder used during pretraining is removed.
+For all downstream tasks, the pre-trained encoder is retained and the decoder removed.
 The Encoder output is processed through a channel-fusion and temporal pooling module (detailed in the following subsection), after which the resulting representation is fed into task-specific prediction heads, whose specifics are provided in the corresponding downstream task sections (see Sec.3.2.2).
@@
 3.2.2 Downstream Tasks
-Table 3: Summary of downstream EMG datasets.
-Dataset #Subj. $f_{s}[\text{Hz}]$ Ch Size [GB] Placement Window size (ms) Pre-Processing Metrics
-Ninapro DB5 [ninapro_db5] 10 200 16 0.2 Forearm 200, 25% overlap $||$ 1000, 25% overlap BP 20–90 Hz $\rightarrow$ NC 50 Hz Accuracy, F1, AUROC
-EPN-612 [benalcazar_emg-epn-612_2020] 612 200 8 5.5 Forearm 200, 25% overlap $||$ 1000, 25% overlap BP 20–90 Hz $\rightarrow$ NC 50 Hz Accuracy, F1, AUROC
-UC Irvine [emg_data_for_gestures_481] 36 200 8 0.32 Forearm 200, 25% overlap $||$ 1000, 25% overlap BP 20–90 Hz $\rightarrow$ NC 50 Hz Accuracy, F1
-Generic Neuromotor Interface [generic_neuromotor_interface_2025] 100 2000 16 31.1 Hand/Arm 8000 (no overlap) HP ($40\text{Hz}$) CLER
-Ninapro DB8 [ninapro_db8] 12 2000 16 23.6 Forearm 200 MAE
-Gaddy Silent Speech [gaddy-klein-2020-digital] 1 1000 8 3.9 Face/Neck $\approx$3000 Series of NC ($60,...,420]\text{Hz} \rightarrow$ HP $2\text{Hz}$ $\rightarrow$ downsampling WER
-BP: band-pass filter; NC: notch filter.
+Table 1: Summary of Downstream Datasets.
+Dataset #S Hz Ch Loc. Metric
+Hand Gesture Classification
+Ninapro DB5 [14] 10 200 16 FA Acc, F1
+EPN-612 [15] 612 200 8 FA Acc, F1
+UCI-EMG [16] 36 200 8 FA Acc, F1
+GNI [5] 100 2k 16 H/A CLER
+Hand Kinematic Regression
+Ninapro DB8 [17] 12 2k 16 FA MAE
+Speech Production & Recognition
+Gaddy [18] 1 1k 8 F/N WER
+Loc: Location (FA: Forearm, H/A: Hand/Arm, F/N: Face/Neck).
+Pre-processing and windowing details are provided in Section III-B.
 The pretrained encoder is evaluated on the datasets listed in Table 1.
@@
 This task consists of classifying different types of hand-wrist movements and is evaluated on the following datasets:
-  • Ninapro DB5 [ninapro_db5] contains data collected from 10 intact subjects, using two Thalmic Myo Armbands (16 channels, $200~\text{Hz}$). The task consists of classifying 52 hand-wrist movements.
-  • EPN-612 [benalcazar_emg-epn-612_2020] contains data collected from 612 subjects, using the Myo Armband (8 channels, $200~\text{Hz}$). The task consists of classifying 5 hand movements.
-  • UC Irvine [emg_data_for_gestures_481] contains data from 36 subjects (8 channels, $200~\text{Hz}$). The task consists of classifying 7 movements.
+  • Ninapro DB5 [39]: data from 10 intact subjects, using two Thalmic Myo Armbands (16 channels, $200~\text{Hz}$). The task consists of classifying 52 hand-wrist movements.
+  • EPN-612 [6] data from 612 subjects, using the Myo Armband (8 channels, $200~\text{Hz}$). The task consists of classifying 5 hand movements.
+  • UC Irvine [32] data from 36 subjects (8 channels, $200~\text{Hz}$). The task consists of classifying 7 movements.
 Datasets are pre-processed using a band-pass filter (Butterworth, $4^{th}$ order, $20-90]~\text{Hz}$), followed by a notch filter ($50~\text{Hz}$).
@@
 We perform an ablation over two different window sizes: $200~\text{ms}$ with 25% overlap, and $1000~\text{ms}$ with 25% overlap.
-After filtering and windowing, each dataset is normalized using per-channel z-score normalization.
-Compared to the pre-training, datasets with fewer than 16 channels are not zero-padded because tokenization is performed on a per-channel basis; thus, having fewer channels simply results in fewer tokens, eliminating the need for padding.
-For Ninapro DB5, we perform a per-repetition split as in [chen_2025_physiowave], each recording contains six repetitions, and for each subject, we assign whole repetitions to a single split so that no subject data overlaps across splits.
-Specifically, repetitions 1, 3, 4, and 6 are used for training, repetition 2 for validation, and repetition 5 for testing.
-The UCI EMG dataset is split by subject index into three disjoint groups: subjects 1–24 are used for training, subjects 25–30 for validation, and subjects 31–36 for testing.
+Finally, each dataset is normalized using per-channel z-score normalization.
+Compared to the pre-training, datasets with fewer than 16 channels are not zero-padded because tokenization is performed on a per-channel basis: having fewer channels simply results in fewer tokens, eliminating the need for padding.
+For Ninapro DB5, we perform a per-repetition split as in [10], each recording contains six repetitions, and for each subject, we assign whole repetitions to a single split so that no subject data overlaps across splits (repetitions 1-3-4-6 are used for training, rep.
+2 for validation, and rep.
+5 for testing).
+The UCI EMG dataset is split by subject index into three disjoint groups: subjects 1–24 for training, subjects 25–30 for validation, and subjects 31–36 for testing.
 Only subjects 11 and 30 performed the ”extended palm” gesture, so they were excluded from all data splits, hence reducing the total number of gestures to 6.
-For EPN-612, we preserve the original per-user organization [benalcazar_emg-epn-612_2020] and derive splits per user.
+For EPN-612, we preserve the original per-user organization [6] and derive splits per user.
 When constructing the testing split, held-out examples per user are used, ensuring that samples from the same user do not appear across train, validation, or test splits.
@@
 For training, we employ the cross-entropy loss computed over the predicted logits and the corresponding ground-truth labels.
-We evaluate performance using Accuracy, F1-score, and AUROC, with the respective definitions given below:
-Accuracy (micro) $=\frac{1}{N}\sum_{i=1}^{N}1(\hat{y_{i}}=y_{i}),$
-Accuracy (macro) $=\frac{1}{K}\sum_{k=1}^{K}\frac{\mathrm{TP}_{k}+\mathrm{TN}_{k}}{\mathrm{TP}_{k}+\mathrm{TN}_{k}+\mathrm{FP}_{k}+\mathrm{FN}_{k}},$
-F1-score (macro) $=\frac{1}{K}\sum_{k=1}^{K}\frac{2\,\mathrm{TP}_{k}}{2\,\mathrm{TP}_{k}+\mathrm{FP}_{k}+\mathrm{FN}_{k}},$
-AUROC (macro) $=\frac{1}{K}\sum_{k=1}^{K}\mathrm{AUROC}_{k}.$
-where $\hat{y}_{i}$ is the predicted class, $y_{i}$ the ground-truth label, and $\mathrm{TP}_{k}$, $\mathrm{FP}_{k}$, $\mathrm{FN}_{k}$ denote true positives, false positives, and false negatives for class $k$.
+We evaluate performance using Accuracy, F1-score, and AUROC.
 We repeat each experiment five times using five distinct random seeds and report the mean performance across runs.
 An additional dataset is also considered as part of this downstream task:
-  • Generic Neuromotor Interface [generic_neuromotor_interface_2025] contains data collected from 4900 participants, recorded with a custom wrist-band developed by Meta Reality Labs (16 channels, $2\text{kHz}$).
-Data from only 100 subjects are publicly available [metagit], randomly selected from the entire cohort.
-Specifically, we tackle the discrete-gesture task presented in the original work, consisting of the classification of 9 fine finger movements.
+  • Generic Neuromotor Interface [26] contains data collected from 4900 participants, recorded with a custom wrist-band developed by Meta Reality Labs (16 channels, $2\text{kHz}$).
+Data from only 100 subjects are publicly available [1], randomly selected from the entire cohort.
+Specifically, we tackle the discrete-gesture task presented in the original work, i.e., classifying 9 fine finger movements.
 We treat this task separately, compared to the three other datasets presented above, owing to substantial methodological differences in data processing and evaluation protocols.
@@
 We employ the Cross-Entropy Loss function, and we conduct 5 runs with 5 seeds values.
-We evaluate the performance with the Classification Error Rate (CLER), defined as the proportion of ground-truth labels for which the matching prediction is incorrect [generic_neuromotor_interface_2025].
+We evaluate the performance with the Classification Error Rate (CLER), defined as the proportion of ground-truth labels for which the matching prediction is incorrect [26].
 The CLER is computed per-gesture and reported as the average across gestures.
@@
 Downstream 2: Hand Kinematic Regression
-This task consists of predicting continuous hand and wrist movement (regression task).
-We employ the Ninapro DB8 [ninapro_db8] dataset.
-It contains EMG data collected from 12 subjects (16 channels, $2~\text{kHz}$), with 3 sessions per subject.
+This task consists of predicting continuous hand and wrist movement (regression).
+We employ the Ninapro DB8 [31] dataset, containing EMG data from 12 subjects (16 channels, $2~\text{kHz}$), with 3 sessions per subject.
 Data were collected with an instrumented glove from ground-truth.
-The task consists of regressing 5 Degrees of Freedom (DoFs), defined as linear combinations of the 18 DoFs of the instrumented glove [krasoulis_2019_effect].
+The task consists of regressing 5 Degrees of Freedom (DoFs), defined as linear combinations of the 18 DoFs of the instrumented glove.
 Each EMG channel is independently standardized using z-score normalization.
 No filtering is applied since the original data is already released filtered ($4\textsuperscript{th}$ order bandpass Butterworth filter, $10-500~\text{Hz}]$.
-The glove data undergo a mapping from 18 Degrees of Freedom (DoF) to 5 Degrees of Actuation (DoA).
+The glove data undergo a mapping from 18 DoFs to 5 Degrees of Actuation.
 Both EMG and glove data are then segmented into fixed-size windows with no overlap in two different settings: $1000\text{ms}$ and $200\text{ms}$.
@@
 We employ the L1 loss function.
-We report the mean absolute error (MAE), root mean squared error (RMSE), and the coefficient of determination $R^{2}$, defined as:
-$\mathrm{MAE}=\frac{1}{N}\sum_{i=1}^{N}\lvert y_{i}-\hat{y}_{i}\rvert$.
-$\mathrm{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}.$
-$R^{2}=1-\frac{\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}.$
-where $y_{i}$ are the true values, $\hat{y}_{i}$ the predicted values, and $\bar{y}=\frac{1}{N}\sum_{i=1}^{N}y_{i}$. We conduct 5 runs, with 5 random seed values, and report the average.
+We report the mean absolute error (MAE), root mean squared error (RMSE), and the coefficient of determination $R^{2}$, averaged over 5 runs with 5 random seed values.
 Downstream 3: Speech Synthesis and Recognition
 We employ the open-vocabulary dataset presented by Gaddy et al.
-[gaddy-klein-2020-digital].
+[16].
 It comprises 8 channels, sampled at $1000~\text{Hz}$, recorded using wet electrodes, placed on the face.
@@
 In the vocalized setting, EMG data were recoded together with a microphone, serving as ground truth.
-For the silent setting, audio recorded during vocalized experiments is aligned to the corresponding silent EMG data using the dynamic time warping algorithm [gaddy-klein-2020-digital].
-The publicly available dataset contains already-aligned data.
-For both tasks, we adopt the same data processing pipeline as in [gaddy2021improved].
-A cascade of notch filters is applied to suppress the first six harmonics of power line frequency ($60,240,...,420]~\text{Hz}$).
-Additionally, a high-pass filter (Butterworth, $2~\text{Hz}$, 3rd order) is applied.
-Data are downsampled (from $1000~\text{Hz}$ to $516.79~\text{Hz}$, matching the processing pipeline in [gaddygit]).
+For the silent setting, audio recorded during vocalized experiments is aligned to the corresponding silent EMG data using the dynamic time warping algorithm [16].
+For both tasks, we adopt the data processing pipeline of [17].
+A cascade of notch filters is applied to suppress the first six harmonics of power line frequency ($60,240,...,420]~\text{Hz}$), followed by a high-pass filter (Butterworth, $2~\text{Hz}$, 3rd order) and downsampling from to $516.79~\text{Hz}$ [51].
 For this dataset, we tackle two sub-tasks: Speech Synthesis and Speech Recognition.
@@
 First, a Transduction model is trained to reconstruct audio target features (Mel-Frequency Cepstral Coefficients, MFCC).
-It consists of 3 residual blocks (as in [gaddy2021improved]) that downsample the input (1600 samples to 200 samples).
-The 200 samples are the input to the pre-trained Transformer Encoder; we add relative positional embeddings, following [gaddy2021improved].
+It consists of 3 residual blocks (as in [17]) that downsample the input (1600 samples to 200 samples).
+The 200 samples are the input to the pre-trained Transformer Encoder; we add relative positional embeddings, following [17].
 A lightweight linear head follows, with an input dimension of $192\times N_{c}$, output dimension of 26 MFCC.
-The model is trained using as loss function the mean of pair-wise MFCC distances ($L$) plus an auxiliary phoneme loss ($L_{phon}$)  [gaddy2021improved].
-In the case of silent EMG data, DTW is used to align the MFCC features predicted from silent EMG to the sequence of target MFCC features obtained from a parallel vocalized recording.
-Predicted MFCC audio features are fed to a HiFi-GAN [kong_2020_hifigangenerativeadversarialnetworks] vocoder, which converts the audio features into raw audio waveforms.
-We adopt the publicly available [gaddygit] pre-trained HiFi-GAN vocoder from Gaddy’s work [gaddy-klein-2020-digital], without any fine-tuning.
-For model evaluation, we employ the Wav2Vec2 [baevski_2020_wav2vec20frameworkselfsupervised] ASR model, pretrained on LibriSpeech (English Language) within SpeechBrain [speechbrain].
+The model is trained using as loss function the mean of pair-wise MFCC distances ($L$) plus an auxiliary phoneme loss ($L_{phon}$)  [17].
+For silent EMG data, DTW is used to align the MFCC features predicted from silent EMG to the sequence of target MFCC features obtained from a parallel vocalized recording.
+Predicted MFCC audio features are fed to a HiFi-GAN vocoder, which converts the audio features into raw audio waveforms.
+We adopt the publicly available pre-trained HiFi-GAN vocoder [51] without any fine-tuning.
+For model evaluation, we employ the Wav2Vec2 [4] ASR model, pretrained on LibriSpeech (English Language) within SpeechBrain [41].
 Speech Recognition consists of recognizing speech (words) without any acoustic output, yielding text as output.
-Producing text directly from EMG signals is preferable to first synthesizing audio and then applying an ASR system: direct text prediction serves as a more efficient surrogate for human intelligibility, avoids the intermediate audio-generation stage, and yields improved recognition accuracy [gaddy_2022_phdthesis].
-Moreover, EMG-to-text models do not require an additional vocoder for speech production, substantially reducing model size and computational load.
+Direct EMG-to-text prediction avoids intermediate audio synthesis and reduces model complexity, yielding improved recognition accuracy [18].
 This makes the approach more suitable for deployment on resource-constrained embedded platforms.
-The model architecture is the same as before (Residuals, pre-trained Transformer Encoder and linear layer with softmax).
+The model architecture is the same as before.
 The input to the final linear layers has a dimension of $N_{c}\times 192$, the output dimension is 37 (26 lowercase English letters, 10 digits, and space character).
-We adopt the Connectionist Temporal Classification (CTC) [graves_2006_connectionist] loss function with a 4-gram language model trained on the LibriSpeech dataset [LibriSpeech].
+We adopt the Connectionist Temporal Classification (CTC) [20] loss function with a 4-gram language model trained on the LibriSpeech dataset [34].
 During inference, beam search is applied to explore the search space and select the paths with the highest likelihood.
-The dataset is split as in Gaddy et al.
-[gaddy2021improved].
+The dataset is split as in [17].
 Training is performed using both silent and vocalized EMG recordings; testing is performed on silent data.
@@
 During training, a size-aware sampler is used to group examples, together producing compact batches and a more efficient training stage.
-For both tasks, we use the Word Error Rate (WER) as the evaluation metric, defined as
-$\text{WER}=\frac{S+D+I}{N},$
-where $S$, $D$, and $I$ denote the number of substitution, deletion, and insertion errors, respectively, and $N$ is the total number of words in the reference transcription.
+For both tasks, we use the Word Error Rate (WER) as the evaluation metric, defined as $\text{WER}=\frac{S+D+I}{N}$, where $S$, $D$, and $I$ denote the number of substitution, deletion, and insertion errors, respectively, and $N$ is the total number of words in the reference transcription.
 3.2.3 Experimental Settings
-All downstream experiments are performed with the same hyperparameters, as per Table 4.
-Table 4: Hyperparameters for model fine-tuning.
-Hyperparameter Value
-Batch size 32
-Peak LR $5\times 10^{-4}$
-Minimum LR $1\times 10^{-5}$
-LR scheduler Cosine
-Optimizer $(\beta_{1},\beta_{2})$ AdamW (0.9, 0.98)(0.9,\ 0.98)
-Weight decay 0.01
-Max training epochs * 50
-Warm-up epochs 5
-Drop-path 0.10
-Layer-wise LR decay 0.90
-Label smoothing 0.10
-LR: learning rate;
-* Early stopping is adopted, with patience of 7 epochs (validation set).
-We evaluate the downstream tasks under two different settings: (i) Full-supervised (FS), where the model is trained from scratch on the task-specific dataset without relying on pre-trained weights; (ii) Full Finetuning (FT), where the entire model is fine-tuned using layer-wise learning rate decay to avoid catastrophic forgetting [kenneweg_forgetting_22].
+All downstream experiments use the AdamW optimizer ($\beta_{1}=0.9,\beta_{2}=0.98$) with a batch size of 32, weight decay 0.01, and label smoothing 0.1.
+We employ a cosine learning rate schedule (50 epochs, 5 warm-up) peaking at $5\times 10^{-4}$ (min $1\times 10^{-5}$) with 0.9 layer-wise decay and 0.1 drop-path.
+Early stopping is adopted with a patience of 7 epochs.
+We evaluate the downstream tasks under two different settings: (i) Full-supervised (FS), where the model is trained from scratch on the task-specific dataset without relying on pre-trained weights; (ii) Full Finetuning (FT), where the entire model is fine-tuned using layer-wise learning rate decay to avoid catastrophic forgetting [28].
 3.3 Model Deployment
-We deploy the TinyMyo on the GreenWaves Technologies GAP9 [greenwaves_gap_sdk] processor, a multi-core RISC-V MCU designed for ultra-low-power edge AI.
+We deploy the TinyMyo on the GreenWaves Technologies GAP9 [21] processor, a multi-core RISC-V MCU designed for ultra-low-power edge AI.
 The architecture features a hierarchical memory organization consisting of a large off-chip L3 (HyperRAM), a 1.5 MB on-chip L2 (shared memory), and a 128 kB L1 (Tightly Coupled Data Memory) shared among a cluster of 8 compute cores and a dedicated cluster controller.
 To enable the execution of the Transformer-based architecture, which exceeds the on-chip L2 capacity in terms of parameters and intermediate activation maps, we developed a custom deployment toolchain that implements a hierarchical streaming architecture.
-Unlike traditional MCU inference engines that require the full model to reside in on-chip memory, our approach treats L2 as a streaming buffer. We implement a multi-level tiling strategy:
+Unlike traditional MCU inference engines that require the full model to reside in on-chip memory, we treat L2 as a streaming buffer. We implement a multi-level tiling strategy:
   • L3 $\rightarrow$ L2: Large tensors (weights and activations) are stored in external L3 memory.
@@
 All intermediate tensors, including those in MHSA (with integer softmax) and GELU/LayerNorm, are stored and processed in INT8 format, with only the LayerNorm scale and shift parameters ($\gamma$ and $\beta$) retained in FP32.
-The only remaining FP32 operation is the final classification layer, which consumes INT8 inputs and weights but accumulates to FP32 logits to enable a direct comparison with the original FP32 baseline; in principle, this layer could also be executed entirely in INT8.
+The only remaining FP32 operation is the final classification layer, which consumes INT8 inputs and weights but accumulates to FP32 logits to enable a direct comparison with the original FP32 baseline; in principle, this layer could also be executed in INT8.
 To maximize the utilization of the limited 1.5 MB L2 memory, we employ an offline liveness analysis pass.
@@
 Tensors are assigned offsets within this arena based on their execution lifetime, allowing mutually exclusive buffers (e.g., non-overlapping intermediate feature maps) to share the same physical memory addresses.
-To maintain numerical fidelity while avoiding costly floating-point operations, all transformer primitives are implemented in fixed-point arithmetic[kim_2021_bert].
+To maintain numerical fidelity while avoiding costly floating-point operations, all transformer primitives are implemented in fixed-point arithmetic[29].
 We use integer-only variants of softmax (i-Softmax), layer normalization (i-LayerNorm), and GELU activation (i-GELU), which leverage lookup tables and multiplicative inverses to approximate their floating-point counterparts with minimal error.
-As proof of concept, we deploy the Hand Gesture Classification downstream task on the NinaPro DB5 dataset.
-Table 5: FLOPs breakdown
-Experiment FLOPs
-Model Pretraining (Encoder + Decoder) 4.001 G
-Ninapro DB5 4.813 G
-EPN-612 1.916 G
-UCI-EMG 1.916 G
-Generic Neuromotor 13.6315 G
-Ninapro DB8 4.852 G
-Speech Production 1.0956 G
-Speech Recognition 1.0921 G
+As proof of concept, we deploy the Hand Gesture Classification downstream task on the EPN-612 dataset.
+4 Results
+4.1 Pre-training
+Figure 2: Example of reconstruction for an EMG window by TinyMyo. Grey areas: masked regions. White areas: unmasked regions.
+Figure 2 shows an example of the reconstruction of the EMG signal performed by the model.
+With a parameter count of only 3.6 million (with only 0.1% of parameters for the decoder), in line with previous works [11], our model demonstrates that high reconstruction fidelity can be maintained even under a strict parameter budget and low computational requirements In terms of computational complexity, the model requires 4.0 GFLOPs for pre-training.
+Inference costs vary by downstream task complexity: approximately 1.1 GFLOPs for speech tasks, 1.9 GFLOPs for EPN-612/UCI-EMG, 4.8 GFLOPs for Ninapro datasets, and 13.6 GFLOPs for the Generic Neuromotor.
 Table 2: Hand Gesture Classification Task — Comparison with State of the Art.
@@
 Acc F1 AUROC Acc F1 AUROC Acc F1 AUROC
-Moment (as in [chen_2025_physiowave]) 385M 86.41 74.42 N.A. 90.87 90.16 N.A. 90.45 91.75 N.A.
+Moment (as in [10]) 385M 86.41 74.42 N.A. 90.87 90.16 N.A. 90.45 91.75 N.A.
 […]
-OTIS (as in [chen_2025_physiowave]) 45M 85.31 72.61 N.A. 87.55 88.03 N.A. 90.62 89.28 N.A.
+OTIS (as in [10]) 45M 85.31 72.61 N.A. 87.55 88.03 N.A. 90.62 89.28 N.A.
 […]
-PhysioWave-Small [chen_2025_physiowave] 5M 84.78 72.54 N.A. 93.12 93.40 N.A. 90.35 89.51 N.A.
+PhysioWave-Small [10] 5M 84.78 72.54 N.A. 93.12 93.40 N.A. 90.35 89.51 N.A.
 […]
@@
 […]
-Notes: PhysioWave [chen_2025_physiowave] use windows of $\approx 1953\text{ms}$.
+Notes: PhysioWave [10] use windows of $\approx 1953\text{ms}$.
 EMG data are upsampled at $2\text{kHz}$ and a fixed window size of 1024 samples is used.
-WaveFormer [chen_2025_waveformer] uses a fixed window of 1024 samples; no information is available on whether resampling is applied.
+WaveFormer [11] uses a fixed window of 1024 samples; no information is available on whether resampling is applied.
 FS: model trained from scratch, FT: fine-tuning pretrained model.
-4 Results
-4.1 Pre-training
-Figure 2: Example of reconstruction for a EMG window by the proposed Foundation model. Grey areas correspond to the masked signal regions, white areas to the unmasked.
-Figure 2 shows an example of the reconstruction of the EMG signal performed by the model.
-With a parameter count of only 3.6 million (with only 0.1% of parameters for the decoder), in line with previous works [chen_2025_waveformer], our model demonstrates that high reconstruction fidelity can be maintained even under a strict parameter budget and low computational requirements A breakdown of the floating-point operations (FLOPs) is available in Table 5.
 4.2 Down-stream Tasks Results
 1. Hand Gesture Classification
-Table 6 reports the results obtained on the Ninapro DB5[ninapro_db5], EPN-612 [benalcazar_emg-epn-612_2020] and the UC Irvine[emg_data_for_gestures_481] datasets, offering a comparison with previous works based on foundation models.
+Table 2 reports the results obtained on the Ninapro DB5, EPN-612, and the UC Irvine datasets, comparing TinyMyo to previous FM works.
 On the NinaPro DB5 dataset, the fine-tuned model trained with a window size of $200\,\text{ms}$ achieves state-of-the-art performance, with an accuracy of $89.41\pm 0.16\,\%$ and an F1-score of $77.97\pm 0.59\,\%$.
@@
 To ensure a fair comparison with prior work, we did not apply any rebalancing techniques during training or evaluation.
-We assume that earlier studies reported the micro accuracy, as no alternative definition was explicitly stated.
-When computing the macro-averaged accuracy score, the accuracy decreases to $75.80\pm 0.62\,\%$, which reflects the imbalance in the dataset.
+Computing a macro-averaged accuracy score reduces the accuracy to $75.80\pm 0.62\,\%$, as expected for an imbalanced dataset.
 On the EPN-612 dataset we achieve SoA results (window size of $1000\text{ms}$, with a global accuracy (FT) of $96.74\pm 0.09\%$, F1-score of $96.74\pm 0.08\%$ and AUROC of $99.70\pm 0.02\%$.
@@
 Model Params CLER Inference Method
-LSTM, Kaifosh [generic_neuromotor_interface_2025] (a) 6.4M 0.1819 Full sequence
-LSTM, Kaifosh [generic_neuromotor_interface_2025] (b) 6.4M 0.1596 Windowed
+LSTM, Kaifosh [26] (a) 6.4M 0.1819 Full sequence
+LSTM, Kaifosh [26] (b) 6.4M 0.1596 Windowed
 […]
 […]
-(a): Full sequence results obtained with methods provided by the authors ([generic_neuromotor_interface_2025]) (b): results obtained with the same model as in a, with the windowed-inference method described in the corresponding Methods section.
+(a): Full sequence results obtained with methods provided by the authors ([26]) (b): results obtained with the same model as in a, with the windowed-inference method described in the corresponding Methods section.
 FS: model trained from scratch, FT: fine-tuning pretrained model
-For the Generic Neuromotor Interface - discrete gesture task [generic_neuromotor_interface_2025], we achieve a CLER of $0.144\pm 0.004$ on the model trained from scratch, which compares to the 0.1596 (obtained with our windowed inference method, see Sec.
+For the Generic Neuromotor Interface - discrete gesture task [26], we achieve a CLER of $0.144\pm 0.004$ on the model trained from scratch, which compares to the 0.1596 (obtained with our windowed inference method, see Sec.
 3.2.2) obtained with the LSTM model proposed in the original work, however with $\approx 44\%$ reduction in model size.
-In this case, the results achieved with the pre-trained Foundation Model are worse ($0.153\pm 0.006$), which we attribute to the intrinsic nature of the problem: to emulate the sequential processing behaviour of the original stacked LSTMs, we imposed a causal mask on the attention mechanism.
+In this case, the results achieved with the pre-trained FM are worse ($0.153\pm 0.006$), which we attribute to the intrinsic nature of the problem: to emulate the sequential processing behaviour of the original stacked LSTMs, we imposed a causal mask on the attention mechanism.
 This constraint restricts our architecture, initially optimized for bidirectional dependency modelling, to a strictly left-to-right context, leading to a degradation in the effectiveness of the pre-trained weights when the future information on which they were originally trained is no longer available.
 Table 3 reports a summary of the results.
-We note that our work is the first to include this dataset among the downstream tasks and to propose an alternative architecture compared to the one employed in the original work.
+We note that our work is the first to include this dataset among the downstream tasks and to propose an alternative architecture compared to the one of the original work.
 2. Hand Kinematic Regression
@@
 Model Params MAE RMSE R2
-TEMPONet TCN, Zanghieri [zanghieri_2021_db8] <<500K 6.89
+TEMPONet TCN, Zanghieri [58] <<500K 6.89
 […]
-DNKF, Bao [BAO202188]∗∗ 13.5
+DNKF, Bao [5]∗∗ 13.5
 […]
@@
 5 DoFs, subject-specific model; ∗∗ 3 DoFs, subject-specific model;
-∗∗∗ 5 DoFss, across-subject model;
-FS: model trained from scratch, FT: fine-tuning pretrained model
+∗∗∗ 5 DoFs, across-subject model; FS: trained from scratch; FT: fine-tuning
 We obtain a MAE of $8.77\pm 0.12^{\circ}$, with a window size of $1000\text{ms}$.
-This value is higher than the result reported by Zanghieri et al.
-[zanghieri_2021_db8] (MAE $6.89^{\circ}$), primarily due to differences in both the training and evaluation protocols.
+This value is higher than the result of.
+[58] (MAE $6.89^{\circ}$), primarily due to differences in both the training and evaluation protocols.
 In our setting, the model is trained in an across-subject regime using data aggregated from all participants, while prior work trained subject-specific models, i.e., one model per individual.
@@
 Method Model Params [M] Model Params [M]
-Transformer, Gaddy [gaddy2021improved] CNN + Transformer 54 HiFi-GAN 13.92 DeepSpeech 36%
-Diff-ETS, Ren [ren_speech_24] as [gaddy2021improved] + Diffusion Probabilistic Model >>54 (a) HiFi-GAN 13.92 Whisper Medium (EN) 32%
-SU-E2S, Scheck. [Scheck2023SUE2S] HuBERT-Soft N.A. HiFi-GAN 13.92 Whisper Medium (EN) 40%
-DiffMV-ETS [Scheck2025DiffMVETSDM] DiffMV-ETS $\approx$ 66 (b) BigVGAN 112 Distil-Whisper-v3 N.A.
-EMGVox-GAN [SUALIHEEN2025101754] as [gaddy2021improved] as [gaddy2021improved] EMGVox-GAN 12 N.A. 36%
-SWS, Lee [lee_speech_2025] as [gaddy2021improved] 54 Sovits-SVC N.A. Whisper Medium (EN) 38%
+Transformer, Gaddy [17] CNN + Transformer 54 HiFi-GAN 13.92 DeepSpeech 36%
+Diff-ETS, Ren [42] as [17] + Diffusion Probabilistic Model >>54 (a) HiFi-GAN 13.92 Whisper Medium (EN) 32%
+SU-E2S, Scheck. [45] HuBERT-Soft N.A. HiFi-GAN 13.92 Whisper Medium (EN) 40%
+DiffMV-ETS [44] DiffMV-ETS $\approx$ 66 (b) BigVGAN 112 Distil-Whisper-v3 N.A.
+EMGVox-GAN [48] as [17] as [17] EMGVox-GAN 12 N.A. 36%
+SWS, Lee [33] as [17] 54 Sovits-SVC N.A. Whisper Medium (EN) 38%
 […]
 […]
-(a) = estimated. The authors employ the same Transduction model as in Gaddy [gaddy2021improved], followed by a Diffusion Model (including residual blocks, attention and convolutional layers)
+(a) = estimated.
+The authors employ the same Transduction model as in Gaddy [17], followed by a Diffusion Model (including residual blocks, attention and convolutional layers).
 (b) = estimated as sum of speaker encoder (6M), prior encoder (30M), and diffusion decoder (30M)
-FS: model trained from scratch, FT: fine-tuning pretrained model
+FS: model trained from scratch. FT: fine-tuning pretrained model
 Table 5 reports the results for the Speech Synthesis Task.
-Our model achieves a WER (FT) of $33.54\pm 1.12\%$, in line with previous works (SoA WER is 32%, achieved by Ren et al.
-[ren_speech_24]).
+We achieve a WER (FT) of $33.54\pm 1.12\%$, in line with previous works (SoA WER is 32%, achieved by Ren et al.
+[42]).
 Compared to previous solutions, our Transduction model offers a significant reduction in parameter size, with 62.5% reduction compared to Scheck et al.
-[Scheck2023SUE2S] and 91.7% reduction compared to Gaddy et al.
-[gaddy2021improved].
-The reported number of parameters accounts only for Transduction model (encoder and model head).
-The number of parameters for the pre-trained Hi-Fi Gan vocoder is 13.92M (same as in [gaddy2021improved, ren_speech_24]).
+[45] and 91.7% reduction compared to Gaddy et al.
+[17].
+The reported number of parameters accounts only for Transduction model (encoder and model head). The number of parameters for the pre-trained Hi-Fi Gan vocoder is 13.92M (same as in [17, 42]).
 Table 6: Silent Speech Recognition Task - Comparison with SoA
 Method Params WER
-Transformer, Gaddy [gaddy_2022_phdthesis] 54 M 28.8%
-MONA, Benster [benster_2024_mona] (a) n.a. 22.2%
-MONA LISA, [benster_2024_mona] (a) n.a. 12.2%
+Transformer, Gaddy [18] 54 M 28.8%
+MONA, Benster [7] (a) n.a. 22.2%
+MONA LISA, [7] (a) n.a. 12.2%
 […]
-(a): Multimodal model
-FT: fine-tuning pretrained model
+(a): Multimodal model. FT: fine-tuning pretrained model
 Table 6 reports the results for the Speech Recognition Task.
@@
 We achieve a WER of $33.95\%\pm 0.97\%$, higher compared to 28.8% reported of Gaddy et al.
-[gaddy_2022_phdthesis], however, with a 91.7% reduction in parameter cont.
-The MONA LISA model [benster_2024_mona] achieves a WER of 12.2%.
+[18], however, with a 91.7% reduction in parameter cont.
+The MONA LISA model [7] achieves a WER of 12.2%.
 However, this performance relies on a multimodal training setup, where both EMG and audio data are jointly used.
@@
 4.3 Model Deployment
-We deploy TinyMyo on the GAP9 microcontroller, as described in Section 3.3.
-The model comprises 3.56 million parameters, with the majority residing in the 8 transformer blocks.
+We deploy TinyMyo on the GAP9 microcontroller (MCU), as described in Section 3.3.
+We achieve an inference time of $0.785~\text{seconds}$ (291.3 M cycles), with energy consumption of $44.90~\text{mJ}$ and average power-envelope of $57.18~\text{mW}$, with GAP9 operating at 370 MHz and using Our custom deployment achieves a high compute utilization of 88.4%; DMA load/store operations account for only $0.4\%$ of cycles, while idle time and overhead are limited to $11.2\%$.
+Crucially, we achieve a DMA/Compute overlap of $99.4\%$, effectively hiding data transfer latencies.
+The model comprises 3.56M parameters, with the majority residing in the 8 transformer blocks.
 Each block contains a multi-head self-attention (MHSA) layer with 3 heads and a feed-forward network (FFN) with an expansion ratio of 4.
-Table 11 reports a computational cost breakdown of the deployed model for one Transformer Block.
-We can observe that the computational cost is dominated by the MHSA QK/AV operations (40% of total MACs) and the MLP projections (40% of total MACs).
-In particular, the QK and AV terms scale as $\mathcal{O}(N^{2}d)$ with the sequence length $N=800$, making them the asymptotic computational bottleneck of the Transformer block.
-With GAP9 operating at 370 MHz, utilizing all 9 cores in the cluster to execute TinyMyo computations in parallel, we achieve an inference in around $12.2\text{seconds}$ at only $0.44\text{J}$ with an average power-envelope of $36.45\text{mW}$.
-Table 11: Computational Cost per Transformer Block
-Component MACs % of Total
-MHSA Q/K/V projections 88M 15%
-MHSA QK scores 123M 20%
-MHSA AV context 123M 20%
-MHSA output projection 29M 5%
-MLP FC1 118M 20%
-MLP FC2 118M 20%
-Notes: QK scores and AV context is (800 $\times$ 800 $\times$ 3 heads). FC1 and FC2 are $192\to 768$ and $768\to 192$ with $4\times$ MLP expansion factor.
+Table 7 summarize the runtime breakdown.
+The overall utilization is high with limited DMA overhead.
+The total cycle count is dominated by the eight Transformer blocks.
+The computational cost for one Transformer block (Table 7) is dominated by the Multi-Head Attention (51.9% of total cycles).
+In particular, the Q/K/V Projections, QK Attention Scores, and AV terms scale as $\mathcal{O}(N^{2}d)$ with the sequence length $N=400$, making them the asymptotic computational bottleneck of the Transformer block.
+Crucially, the INT8 kernels reach strong throughput, with several attention and MLP sub-operations exceeding 991010 MACs/cycle (Table 7 The MLP blocks in the model head account for 33.7% of the total cycles.
+)
+Table 7: Sub-operation Breakdown for one transformer block
+Operation Cycles (M) % MACs (M) MACs/Cyc
+Pre-Attention Norm 2.2 6.2% 0.4 0.17
+Multi-Head Attention 18.6 51.9% 120.4 6.46
+   Q/K/V Projections 5.7 15.9% 44.2 7.76
+   Permute/Reshape 0.1 <<0.1%
+   QKT Attention Scores 3.4 9.4% 30.7 9.08
+   Softmax Normalization 4.5 12.6%
+[…]
+   Output Projection 1.9 5.3% 14.7 7.75
+Pre-MLP Norm 2.1 5.9% 0.4 0.18
+[…]
+GELU Activation 0.8 2.3% 3.1 3.73
+[…]
+Total 35.9 100% 242.2 6.74
 5 Conclusion
-In this work, we presented TinyMyo, the first compact EMG FM capable of broad generalization across diverse downstream tasks and real-time edge deployment on an ultra-low-power microcontroller.
+We presented TinyMyo, the first compact EMG FM capable of broad generalization across diverse downstream tasks and real-time edge deployment on an ultra-low-power MCU.
 TinyMyo is based on a Transformer-encoder architecture, with a size of 3.6 million parameters.
@@
 For gesture recognition, we obtain SoA accuracy on the NinaPro DB5 ($89.41\pm 0.16\,\%$), EPN-612 ($96.74\pm 0.09\,\%$), and UCI-EMG ($97.56\pm 0.32\,\%$) datasets.
-Furthermore, on the Generic Neuromotor Interface – discrete-gesture dataset, TinyMyo achieves a CLER of $0.153\pm 0.006$ with an approximate $44\,\%$ reduction in model size compared to the original work.
+Furthermore, on the Generic Neuromotor Interface – discrete-gesture dataset, TinyMyo achieves a CLER of $0.153\pm 0.006$ with an approximate $44\,\%$ reduction in model size compared to SoA.
 On the hand kinematic regression task, we achieve a MAE of $8.77\pm 0.12$ with a subject-agnostic model.
@@
 Finally, we demonstrate that TinyMyo can also adapt to the speech recognition task, achieving a WER of $33.95\pm 0.97\%$.
-TinyMyo is also the first FM for EMG deployed on the edge, achieving an inference in $12.2\text{s}$ at $0.44\text{J}$, with an average power-envelope of $36.45\text{mW}$.
-We identify the following limitations, which will be addressed as part of future work:
+TinyMyo is also the first FM for EMG deployed on the edge, running in only $0.785~\text{s}$, with an energy of $44.91~\text{mJ}$ and an average power-envelope of $57.18~\text{mW}$.
+Future work will focus on the following limitations:
   • The datasets used for pre-training consist primarily of EMG recordings from the forearm and wrist during gesture-related tasks.
@@
   • Further studies are required to assess the model’s generalization capabilities in zero-shot settings and to quantify the amount of labeled data needed to achieve strong performance on novel downstream tasks.
-  • Despite demonstrating, to the best of our knowledge, the first deployment of an FM for EMG processing on an ultra–low-power microcontroller, the inference time remains too high for seamless use in real-world applications.
-Future work should therefore focus on optimizing the network architecture to minimize the computation needed for inference.
-This could be achieved by doing an ablation study on reducing the sequence length, by adopting a cheaper attention mechanism [dimofte_2025_cerebro] or exploring more aggressive quantization techniques (e.g.: sub 8-bit [dong_2023_packqvit].
-More efficient architectures, such as state-space models (e.g., MAMBA [gu_2024_mamba], FEMBA [tegon_2025_femba]), whose computational cost scales linearly with the context length, would also be worth exploring.
-Moreover, distributing the computation over multiple GAP9 nodes or deploying it on more powerful platforms such as Siracusa [prasad_2024_siracusa] is expected to yield substantial speed-ups.
-Architectural improvements and refinements in the training pipeline also represent an important future direction.
-We aim to explore alternative pooling mechanisms (e.g., attention-based) to weight per-channel contribution more precisely.
-By open-sourcing111after review our pre-trained models and the architectures used for downstream evaluations, we aim to provide a versatile resource that can accelerate future research and serve as a foundation for the broader EMG community.
-Our intention is that other researchers will extend this backbone with additional datasets, improved training strategies, and architectural variations, aligning with the community-wide effort recently advocated by Eddy et al.
-[eddy_2025_raising].
+  • Despite demonstrating the first deployment of an FM for EMG on an ultra–low-power MCU, the inference time is still too high for full real-time operation in low-latency closed loop applications.
+Future work will focus on optimizing the network architecture to minimize the computation needed for inference.
+This could be achieved by doing an ablation study on reducing the sequence length, by adopting a cheaper attention mechanism [12], by exploring more aggressive quantization techniques (e.g.: sub 8-bit [14], and by evaluating state-space models [22, 49], whose computational cost scales linearly with the context length.
+Also, distributing the computation over multiple GAP9 nodes or deploying it on more powerful platforms [40] would yield substantial speed-ups.
+By open-sourcing [^1] https://github.com/pulp-bio/BioFoundation our pre-trained models and the architectures used for downstream evaluations, we aim to provide a versatile resource that can accelerate future research and serve as a foundation for the broader EMG community.
+Our intention is that other researchers will extend this backbone with additional datasets, improved training strategies, and architectural variations.
 Acknowledgment
-The authors thank Chen Yanlong (ETH Zürich) for technical support and Lorenzo Lamberti (ETH Zürich) for fruitful discussions.
+We thank Chen Yanlong (ETH Zürich) for technical support and Lorenzo Lamberti (ETH Zürich) for fruitful discussions.